### v1

# Higher-order homogenised riblet boundary conditions

Abstract
The description of riblets and other drag-reducing devices has long used the concept of longitudinal and transverse protrusion heights, both as a means to predict the drag reduction itself and as equivalent boundary conditions to simplify numerical simulations by transferring the effect of riblets onto a flat virtual boundary. The limitation of this idea is that it stems from a first-order approximation in the riblet-size parameter $s^{+}$, and as a consequence it cannot predict other than a linear dependence of drag reduction upon $s^{+}$; in other words, the initial slope of the drag-reduction curve. Here the concept is extended to a full asymptotic expansion using matched asymptotics, which consistently provides higher-order protrusion heights and higher-order equivalent boundary conditions on a virtual flat surface. This procedure also allows us to explore nonlinearities of the Navier-Stokes equations and the way they enter the $s^{+}$-expansion, with somewhat surprising results.
1 Introduction
Surfaces that are not smooth alter the flow-induced transfer of momentum, heat and mass (Jiménez, 2004; Bottaro, 2019; García-Mayoral et al., 2019; Chung et al., 2021). Examples include riblets (García-Mayoral & Jiménez, 2011a), shark denticles (Lang et al., 2008; Savino & Wu, 2024), seal fur (Itoh et al., 2006; de Segura & García-Mayoral, 2019), superhydrophobic or liquid-infused surfaces (García-Cartagena et al., 2019), roughness (Busse et al., 2017; Thakkar et al., 2018; Flack & Schultz, 2010, 2014) and wavy surfaces (Charru et al., 2013; Thorsness et al., 1978). Understanding the relationship between form (shape, topography, morphology) and function (alteration of transfer properties) of these surfaces is not trivial, generally requiring challenging laboratory experiments (Abrams & Hanratty, 1985; Bechert et al., 1997; Schultz & Flack, 2013) or numerical simulations (Busse et al., 2017; García-Mayoral & Jiménez, 2011b). However, in regimes where a small parameter exists, inroads can be made with asymptotic analyses.
As discussed in Luchini (2013), surface texture characterized by a small scale can be analysed in two separate distinguished limits, one of which is a shallow-roughness limit where height of the texture tends to zero while its horizontal scales stay fixed. These are shallow or wavy surfaces (Luchini & Charru, 2019; Abrams & Hanratty, 1985). The other limit is the small-roughness limit where all three dimensions tend to zero proportionally to each other, i.e. the texture stays geometrically self-similar, e.g. riblets (García-Mayoral & Jiménez, 2011b; Endrikat et al., 2021; Modesti et al., 2021) and rough surfaces (Thakkar et al., 2018; Sharma & García-Mayoral, 2020) with small surface features compared to flow features. In this paper we shall be concerned with the small-roughness limit. In particular, we will develop a higher-order matched asymptotic expansion that encapsulates the relationship between flow alteration and surface details via equivalent (homogenised) boundary conditions. Apart from their direct use for informing design (e.g. García-Mayoral & Jiménez, 2011a), such boundary conditions circumvent the need to resolve the small surface features in flow simulations, easing a bottleneck that enables routine computations and subsequent analysis towards overall drag savings, for example (Lācis et al., 2020; Ahmed & Bottaro, 2025).
In the context of riblets, Bechert & Bartenwerfer (1989) first developed the idea of the longitudinal protrusion height, which can also be interpreted as a slip length (cf. Bottaro, 2019). When applied in the turbulent-flow regime, it predicts the wall-normal distance below a reference plane (often taken as the plane of the riblet tips) where the mean (longitudinal) flow extrapolates to zero. In this sense, the theory predicts that the mean flow is shifted by riblets, an idea which has been confirmed in numerical simulations (García-Mayoral & Jiménez, 2011b). In the presence of three-dimensional turbulence, it was also necessary to consider the other (non-longitudinal) flow components. This was explored by Luchini et al. (1991), who identified the shift in turbulence to be the lateral- or cross-flow (spanwise to the mean) protrusion height. Then, only can the drag change that modifies the inner layer of wall turbulence be explained and extrapolated to full scale (Luchini, 1996). Specifically, the drag change is proportional to the difference between the longitudinal and cross-flow protrusion heights, which is independent of the reference plane from which the protrusion heights are measured.
The protrusion heights describe well the asymptotic change in drag for vanishingly small roughness in turbulence flow (Wong et al., 2024), however, the effect of wall transpiration, i.e. wall-normal component of motion, needs to also be accounted for where small roughness features are no longer vanishingly small (Ibrahim et al., 2021; Bottaro, 2019). For riblets that are comparable in size to the quasi-streamwise vortices of turbulence, a physically based so-called viscous vortex model was developed by Wong et al. (2024) to allow for this wall-normal component. Another approach is the transpiration-resistance model of Lācis et al. (2020); Khorasani et al. (2022), which are essentially homogenised boundary conditions for the macroscopic flow, with coefficients obtained from microscopic flow calculations performed in a single repeating unit of surface texture. The success of this approach relies on the linearity between the macrosopic flow and the microscopic response, as allowed by Stokes flow. A similar approach is also pursued by Zampogna et al. (2019). To extend to inertial effects within the texture, Zampogna & Bottaro (2016) added an Oseen-like approximation of the advection term, which retains the linearity between the macroscopic forcing the microscopic response. Ahmed & Bottaro (2025) demonstrated the utility of these boundary conditions in direct numerical simulations of turbulence.
A formal higher-order matched-asymptotic expansion of the wall-boundary condition of a periodic surface texture has not been attempted. Although the effect of wall transpiration has been recognised as a second-order effect and modelled to varying degrees (Lācis et al., 2020; Ibrahim et al., 2021; Wong et al., 2024; Ahmed & Bottaro, 2025), it is unclear whether other effects that have been omitted are of the same order. This is one of the questions this paper will address. Another way to view the effect of wall transpiration is the influence of boundary conditions that vary in the directions parallel to the wall. The first-order asymptotics as developed by Luchini et al. (1991) effectively assumed that the boundary conditions are spatially uniform, and so the wall-normal velocity must be zero at this level of approximation for an impermeable wall at the macroscopic flow scale. However, wall-parallel spatial variations in boundary conditions will allow for effective wall transpiration at the flow scale (Lācis et al., 2020; Wong et al., 2024; Hao & García-Mayoral, 2025). Further, we will also see that order-by-order matching in physical space does not require linearity between the macroscopic forcing and the microscopic response, with no ad-hoc modelling nor tuning parameters.
The exercise recovers the first-order protrusion heights (Bechert & Bartenwerfer, 1989; Luchini et al., 1991) and then extends them to higher orders, which contain nonlinear terms that reflect the nonlinearity of Navier–Stokes equations. The general idea that made this possible is to apply matched inner and outer asymptotic expansions in the classical sense of van Dyke (1964) but then, instead of solving the inner and outer equations in a back-and-forth cascade, represent the outer solution by a formal Cauchy series and only (numerically) solve the inner equations. This allows an equivalent boundary condition to be formulated, at any desired order of accuracy, that completely replaces the inner dynamics by its projection on a flat virtual wall. First we start with the 2D Laplace equation to demonstrate the mechanics of the process, and then apply it with increasing complexity to Stokes and then the 3D Navier–Stokes equations. The main operational results, at a first reading that skips the mathematics involved, are collected in the third-order equivalent boundary condition (7.1) and the table of its coefficients for five typical wall textures, table 1.
2 2D Laplace equation
The Laplace equation may serve as an example before handling more complex flow problems.
Figure 1: Setup: $s^{*}$ is the riblet period/spacing, $L^{*}\gg s^{*}$ is the upper boundary condition plane (measured from the tips at $z^{*}=0$) where flow on length scales O$(L^{*})$ is imposed.
Let us consider the solution of the 2D Laplace equation
$\Delta_{2}^{*}\phi^{*}=\frac{\partial^{2}\phi^{*}}{\partial y^{*2}}+\frac{\partial^{2}\phi^{*}}{\partial z^{*2}}=0$ (2.1)
in a 2D strip characterized by a periodic arrangement of protruberances or cavities (riblets) of period $s^{*}$ on one wall (figure 1) and an opposing infinite flat wall at far-away distance $L^{*}\gg s^{*}$. Coordinates are named $y^{*}$ along the wall and $z^{*}$ normal to the wall with an eye to later three-dimensional applications. A superscript is used to mark these quantities as dimensional, as opposed to plain letters used further down to denote dimensionless variables. Also, later this will become a flow problem, but for the time being let the unknown $\phi^{*}$ be any scalar quantity that obeys the Laplace equation, say a temperature field. For definiteness let us also impose a Dirichlet boundary condition ($\phi^{*}=0$) on the $s^{*}$-periodic riblet wall, located near $z^{*}=0$, and a constant Neumann boundary condition ($\phi^{*}_{z^{*}}={}$const.) on the flat wall located at $z^{*}=L^{*}$.
2.1 Wavenumber domain (constant outer conditions)
The solution to the Laplace equation in this geometry is itself periodic and can be represented as a Fourier series
$\phi^{*}(y^{*},z^{*})=\sum_{n=-\infty}^{\infty}\hat{\phi}^{*}_{n}(z^{*})\exp(\mathrm{i}n\beta^{*}_{1}y^{*}),$ (2.2)
where $\beta^{*}_{1}=2\pi/s^{*}$. Far from the periodic wall the mean ($n=0$) component is given in (4) of Luchini et al. (1991),
$\hat{\phi}^{*}_{0}=a^{*}_{0}+b^{*}_{0}z^{*}=\hat{\phi}^{*}_{0,z}(L^{*})(z^{*}+h^{*}),$ (2.3)
whereas all the $n\neq 0$ components decay exponentially as $\exp(-\beta^{*}_{1}z^{*})$ or faster. It follows that in most of the domain (precisely for $\beta^{*}_{1}z^{*}\gg 1$, or $z^{*}\gg s^{*}$) the solution has a linear profile as if the wall were flat, and appears to cross zero at a virtual origin $z^{*}=-h^{*}$. This virtual origin, or its distance from an arbitrary reference location which often is taken to be the riblet tip, by definition is the protrusion height, so named by Bechert & Bartenwerfer (1989) who first introduced and calculated it for the Laplace equation applied to longitudinal flow past riblets. Since the governing equation is linear, and the whole solution $\phi^{*}$ is proportional to the strength of the imposed flux $\hat{\phi}^{*}_{0,z^{*}}(L^{*})$, the protrusion height is independent of $\hat{\phi}^{*}_{0,z^{*}}(L^{*})$ and can be obtained once and for all for a given riblet geometry from a single numerical computation, several examples of which were given in Luchini et al. (1991). In addition, given the scale invariance of the Laplace equation, $h^{*}$ is proportional to $s^{*}$ and can be written as $h^{*}=h_{1}s^{*}$, where $h_{1}$ is a dimensionless number calculated for period equal to unity.
It is also apparent, since the $n\neq 0$ Fourier components decay exponentially with distance and thus also with the ratio $L^{*}/s^{*}$, that the approximation provided by substituting the linear profile (2.3) for the actual solution (2.2) is exponentially accurate (its error decays faster that any finite power of $s^{*}/L^{*}$), and a power expansion in the small parameter $\varepsilon=s^{*}/L^{*}$ at fixed $z^{*}/L^{*}$, such as will be considered in the next section, contains no higher powers of $\varepsilon$ than the first-order term with coefficient $h_{1}$ already included in (2.3).
2.2 Physical domain (slowly varying outer conditions)
Let us now consider a situation where the outer imposed flux is not exactly constant with $y^{*}$ but slowly varying on a length scale $L^{*}\gg s^{*}$. Under these conditions the solution of the Laplace equation will no longer be exactly periodic, and its expansion in powers of $s^{*}/L^{*}$ will contain non-trivial higher-order corrections. (Equivalently the period $s^{*}$ itself could be assumed to be slowly varying, but we shall not consider this extension at present.) When the main field is characterized by a length scale $L^{*}$ and the wall texture by a period $s^{*}$, the problem lends itself to matched asymptotic expansions in the parameter $\varepsilon=s^{*}/L^{*}$. Notice that, even if the Laplace equation is linear with respect to forcing, it is not linear with respect to geometry, and therefore the expansion is allowed to contain higher powers of $\varepsilon$.
In order to introduce matched asymptotic expansions, in the style in which they are usually introduced in boundary layers (van Dyke, 1964), we need inner coordinates $Y,Z$, made dimensionless with the riblet period $s^{*}$, and outer coordinates $y=y_{0}+\varepsilon Y,z=\varepsilon Z$ made dimensionless with the outer scale $L^{*}$ in a neighbourhood of a generic boundary point $(y_{0},0)$. The, now dimensionless, dependent variable $\phi$ can be expanded in an inner series
$\phi=\varPhi_{0}(Y,Z)+\varepsilon\varPhi_{1}(Y,Z)+\cdots=\sum_{i=0}^{\infty}\varepsilon^{i}\varPhi_{i}(Y,Z),$ (2.4)
and an outer series
$\phi=\phi_{0}(y,z)+\varepsilon\phi_{1}(y,z)+\cdots=\sum_{i=0}^{\infty}\varepsilon^{i}\phi_{i}(y,z).$ (2.5)
In consequence of the Laplace equation’s being linear and scale-invariant, each term of both the inner and outer expansions will independently obey its own Laplace equation:
$\Delta_{2}\varPhi_{i}(Y,Z)=0,\quad\Delta_{2}\phi_{i}(y,z)=0\quad\forall i.$
In addition, $\varPhi_{i}$ will obey a Dirichlet boundary condition on a curve $Z=F(Y)$, $\varPhi_{i}[Y,F(Y)]=0$, with the riblet shape $F(Y)$ periodic of period $1$ (because $Y$ has been made dimensionless with the period), and $\phi_{i}$ a variable Neumann boundary condition at the outer wall $z=1$:
$\phi_{0,z}(y,1)=q(y);\quad\ \phi_{i,z}(y,1)=0\;\text{for }i\neq 0.$ (2.6)
In order to match the inner to the outer solution in a neighbourhood of a generic position $y_{0}$, we can set $y=y_{0}+\varepsilon Y$, $z=\varepsilon Z$, and expand each term of the outer series (2.5) in a 2D Taylor series in $\varepsilon Y$ and $\varepsilon Z$:
$\begin{split}\phi={}&\phi_{0}(y_{0},0)+\varepsilon Y\frac{\partial\phi_{0}}{\partial y}(y_{0},0)+\varepsilon Z\frac{\partial\phi_{0}}{\partial z}(y_{0},0)+\varepsilon\phi_{1}(y_{0},0)+{}\\ &\varepsilon^{2}\frac{Y^{2}}{2}\frac{\partial^{2}\phi_{0}}{\partial y^{2}}(y_{0},0)+\varepsilon^{2}YZ\frac{\partial^{2}\phi_{0}}{\partial y\partial z}(y_{0},0)+\varepsilon^{2}\frac{Z^{2}}{2}\frac{\partial^{2}\phi_{0}}{\partial z^{2}}(y_{0},0)+{}\\ &\varepsilon^{2}Y\frac{\partial\phi_{1}}{\partial y}(y_{0},0)+\varepsilon^{2}Z\frac{\partial\phi_{1}}{\partial z}(y_{0},0)+\varepsilon^{2}\phi_{2}(y_{0},0)+{}\cdots.\end{split}$ (2.7)
According to the principle of matched asymptotic expansions, as applied to the small-roughness limit, the inner and outer expansions will not be uniformly valid but only match in an intermediate distinguished limit, in which $Y$ and $Z$ tend to infinity at the same time as $z\rightarrow 0$ and $y\rightarrow y_{0}$ with $\varepsilon\rightarrow 0$. In this limit we shall term by term equate coefficients of equal powers of $\varepsilon$ between (2.4) and (2.7). So doing yields the following hierarchy:
2.2.1 Order zero
At inner order 0, matching yields that $\varPhi_{0,Z}(Y,\infty)=0$, because the $\varepsilon^{0}$ term of (2.7) is just a constant and has zero derivative; being a solution of the Laplace equation with all zero boundary conditions, $\varPhi_{0}$ is identically null (and the inner expansion could just as well have started from $i=1$).
At outer order 0 we need to solve the Laplace equation for $\phi_{0}(y,z)$ with the boundary condition (2.6) at the outer wall and $\phi_{0}(y,0)=0$ at the flat inner wall $z=0$, to which the periodic wall reduces for $\varepsilon=0$. From the result we can extract $\phi_{0,z}(y,0)$ (which also includes information on all of its $y$-derivatives) for later use.
2.2.2 First order
At inner order $1$ from (2.7) we obtain
$\varPhi_{1}(y_{0};Y,Z)\sim Z\phi_{0,z}(y_{0},0)+\phi_{1}(y_{0},0)$ (2.8)
where $y_{0}$ is to be treated as a parameter as far as the Laplace equation for $\varPhi_{1}$ is concerned, and symbol $\sim$ is here precisely defined to mean that the difference between its l.h.s. and r.h.s. must tend to zero when $Z\rightarrow\infty$. This definition may imply more than one actual equation, since the l.h.s. and r.h.s. are asymptotically polynomials in $Z$, and each coefficient of these polynomials must individually match for the difference to tend to zero. Notice also that the second term of (2.7), $Y\phi_{0,y}(y_{0},0)$, has been omitted from (2.8) because $\phi_{0}(y,0)=0$ for all $y$, and its derivatives are thus also zero. Asymptotic condition (2.8) can be split into the two equations
$\varPhi_{1,Z}(y_{0};Y,\infty)=\phi_{0,z}(y_{0},0)$ (2.9)
and
$\phi_{1}(y_{0},0)=\lim_{Z\rightarrow\infty}\left[\varPhi_{1}(y_{0};Y,Z)-Z\phi_{0,z}(y_{0},0)\right]$ (2.10)
Equation (2.9) (the $r.h.s.$ of which will be a non-trivial function of $y_{0}$ unless, as in the previous section, $\phi_{0,z}(y,1)$ is constant) closes the Laplace problem for the order-1 inner solution $\varPhi_{1}$. Equation (2.10) in turn provides the wall boundary condition for the first-order outer solution $\phi_{1}$. Since $\varPhi_{1}\propto\phi_{0,z}(y_{0},0)$ (because of (2.9) and the linearity of the Laplace problem), (2.10) can also be written as
$\phi_{1}(y_{0},0)=h_{1}\phi_{0,z}(y_{0},0)$ (2.11)
where constant $h_{1}$, independent of $y_{0}$, is the same protrusion height as already introduced in §2.1, and is the only quantity in (2.11) which depends on the actual wall texture $Z=F(Y)$. In fact, introducing the normalized $\overline{\varPhi_{11}}=\varPhi_{1}(y_{0};Y,Z)/\phi_{0,z}(y_{0},0)$, one realizes that $\overline{\varPhi_{11}}$ is independent of $y_{0}$, and obeys the Laplace equation $\Delta_{2}\overline{\varPhi_{11}}=0$ with the normalized boundary conditions
$\overline{\varPhi_{11}}[Y,F(Y)]=0,\quad\overline{\varPhi_{11}}_{,Z}(Y,\infty)=1,$ (2.12)
Substituting $\overline{\varPhi_{11}}(Y,Z)$ in (2.10) and then comparing with (2.11), one can obtain
$h_{1}=\lim_{Z\rightarrow\infty}\left[\overline{\varPhi_{11}}(Y,Z)-Z\right].$ (2.13)
Whereas for theoretical purposes the inner-outer expansion is the preferred approach, in numerical computation to have to deal with the expansion is inconvenient. We can, within the same order of approximation, sum the zeroth- and first-order boundary conditions to generate
$\phi(y,0)=\phi_{0}(y,0)+\varepsilon\phi_{1}(y,0)=\varepsilon h_{1}\phi_{0,z}(y,0)+\text{O}(\varepsilon^{2}),$
and, within a similar error O$(\varepsilon^{2})$, replace $\varepsilon\phi_{0,z}(y,0)$ by $\varepsilon\phi_{z}(y,0)$ (similarly, in a sense, to what in Aerodynamics is called “interactive boundary layer theory”, e.g. Cebeci (1999)). There results a first-order equivalent boundary condition for the total $\phi$ without any more reference to its series expansion:
$\phi(y,0)=\varepsilon h_{1}\phi_{z}(y,0)+\text{O}(\varepsilon^{2}),$ (2.14)
or in dimensional form:
$\phi^{*}(y^{*},0)\simeq s^{*}h_{1}\phi^{*}_{z^{*}}(y^{*},0),$ (2.15)
This is the slip-length boundary condition often encountered in the literature on the computational solution of the outer equation, with the slip length $h^{*}=s^{*}h_{1}$ being just another name for the protrusion height. But in addition, as will now be seen, the present matched asymptotic expansion allows the analysis to be continued to higher and higher orders.
2.2.3 Second order
Matching the second-order terms of (2.4) and (2.7) gives:
$\varPhi_{2}(y_{0};Y,Z)\sim YZ\phi_{0,yz}(y_{0},0)+Y\phi_{1,y}(y_{0},0)+Z\phi_{1,z}(y_{0},0)+\phi_{2}(y_{0},0)$ (2.16)
where $\phi_{0,yy}$ has been omitted because $\phi_{0}(y,0)$ is identically zero, and $\phi_{0,zz}$ has also been omitted for the same reason, because from the Laplace equation it equals $-\phi_{0,yy}$. We notice also that, from the $y$-derivative of (2.11), $\phi_{1,y}(y_{0},0)$ can be replaced by $h_{1}\phi_{0,yz}(y_{0},0)$.
A solution of the Laplace equation for $\varPhi_{2}$ such that it grows linearly in $Y$ at large $Y$, like (2.16) requires, can be written as
$\varPhi_{2}=\varPhi_{21}(y_{0};Y,Z)+Y\,\varPhi_{211}(y_{0};Y,Z),$ (2.17)
where $\varPhi_{21}$ and $\varPhi_{211}$ are periodic functions of $Y$ (of period $1$) obeying
$\Delta_{2}\varPhi_{211}(y_{0};Y,Z)=0;\quad\Delta_{2}\varPhi_{21}(y_{0};Y,Z)=-2\varPhi_{211,Y}(y_{0};Y,Z)$ (2.18)
and
$\varPhi_{211,Z}(y_{0};Y,\infty)=\phi_{0,yz}(y_{0},0),$ (2.19)
$\varPhi_{21,Z}(y_{0};Y,\infty)=\phi_{1,z}(y_{0},0).$ (2.20)
This is the second-order inner solution. Owing to linearity of the Laplace equation with respect to its inhomogeneous boundary condition, $\varPhi_{211}$ can be written in normalized form as $\varPhi_{211}(y_{0};Y,Z)=\overline{\varPhi_{211}}(Y,Z)\,\phi_{0,yz}(y_{0},0)$. But $\overline{\varPhi_{211}}$ turns out to obey the same equation and boundary conditions as $\overline{\varPhi_{11}}$, and thus coincides with it. In addition, using (2.11) to replace $\phi_{1,y}$, and $\overline{\varPhi_{211}}=\overline{\varPhi_{11}}\sim Z+h_{1}$ from (2.13), it can be shown that $\varPhi_{211}$ cancels out on both sides of (2.16) as follows:
$\begin{split}&\lim_{Z\rightarrow\infty}\left[Y\varPhi_{211}(y_{0};Y,Z)-YZ\phi_{0,yz}(y_{0},0)-Y\phi_{1,y}(y_{0},0)\right]=\\ &\lim_{Z\rightarrow\infty}\left[Y(Z+h_{1})\phi_{0,yz}(y_{0},0)-YZ\phi_{0,yz}(y_{0},0)-Yh_{1}\phi_{0,yz}(y_{0},0)\right]=0.\end{split}$
The completion of (2.16) then gives
$\phi_{2}(y_{0},0)=\lim_{Z\rightarrow\infty}\left[\varPhi_{21}(y_{0};Y,Z)-Z\phi_{1,z}(y_{0},0)\right]$ (2.21)
as the wall boundary condition for the second-order outer solution.
$\varPhi_{21}$, on the other hand, is comprised of two normalized contributions, one from the r.h.s. of (2.18) and one from the boundary condition (2.20), $\varPhi_{21}=\overline{\varPhi_{21}}\phi_{0,yz}(y_{0},0)+\overline{\varPhi_{212}}\phi_{1,z}(y_{0},0)$. It turns out that $\overline{\varPhi_{212}}$ again coincides with $\overline{\varPhi_{11}}$, whereas $\overline{\varPhi_{21}}$ is the solution of
$\Delta_{2}\overline{\varPhi_{21}}=-2\overline{\varPhi_{11}}_{,Y}$ (2.22)
with homogeneous boundary conditions $\overline{\varPhi_{21}}[Y,F(Y)]=0$, $\overline{\varPhi_{21}}_{,Z}(Y,\infty)=0$. Upon defining a new constant
$h_{2}=\lim_{Z\rightarrow\infty}\overline{\varPhi_{21}}(Y,Z),$ (2.23)
the second-order outer boundary condition (2.21) can be rewritten as
$\phi_{2}(y,0)=h_{1}\phi_{1,z}(y,0)+h_{2}\phi_{0,yz}(y,0),$ (2.24)
and the compound equivalent boundary condition as
$\phi(y,0)=\varepsilon h_{1}\phi_{z}(y,0)+\varepsilon^{2}h_{2}\phi_{yz}(y,0),$ (2.25)
or in dimensional form as
$\phi^{*}(y^{*},0)=s^{*}h_{1}\phi^{*}_{z^{*}}(y^{*},0)+s^{*2}h_{2}\phi^{*}_{y^{*}z^{*}}(y^{*},0).$ (2.26)
The new dimensionless coefficient $h_{2}$ will be called “second-order protrusion height” for naming consistency, but it should be evident that $s^{*2}h_{2}$ is an area and not a height. Also, it can be noted that if $\overline{\varPhi_{21}}$ happens to be an odd function of $Y$, a limit such as (2.23), which must then equal its opposite under the transformation $Y\leftarrow-Y$, is necessarily zero. More about this symmetry at the end of §2.2.5.
Now a pattern begins to appear.
2.2.4 Third order
The third-order matching condition is
$\varPhi_{3}(y_{0};Y,Z)\sim\left(\frac{Y^{2}Z}{2}-\frac{Z^{3}}{6}\right)\phi_{0,yyz}+\frac{Y^{2}-Z^{2}}{2}\phi_{1,yy}+YZ\phi_{1,yz}+Y\phi_{2,y}+Z\phi_{2,z}+\phi_{3},$
where second and higher $z$-derivatives of $\phi_{0}$ and $\phi_{1}$ have been substituted from the Laplace equation, and subsequently $y$-derivatives of $\phi_{0}$ have been cancelled. $\phi_{1,yy}$ and $\phi_{2,y}$ can be replaced from (2.10) and (2.21) as
$\phi_{1,yy}(y_{0},0)=h_{1}\phi_{0,yyz}(y_{0},0)$
$\phi_{2,y}(y_{0},0)=h_{1}\phi_{1,yz}(y_{0},0)+h_{2}\phi_{0,yyz}(y_{0},0),$
to yield
$\varPhi_{3}(y_{0};Y,Z)\sim\left[\frac{Y^{2}}{2}(Z+h_{1})+Yh_{2}-\frac{Z^{3}}{6}-\frac{Z^{2}}{2}h_{1}\right]\phi_{0,yyz}+Y\left(Z+h_{1}\right)\phi_{1,yz}+Z\phi_{2,z}+\phi_{3},$ (2.27)
By analogy with (2.17) one can make the following ansatz:
$\begin{split}\varPhi_{3}=\left[\overline{\varPhi_{31}}(Y,Z)+Y\,\overline{\varPhi_{21}}(Y,Z)+(Y^{2}/2)\,\overline{\varPhi_{11}}(Y,Z)\right]\phi_{0,yyz}+{}\\ \left[\,\overline{\varPhi_{21}}(Y,Z)+Y\overline{\varPhi_{11}}(Y,Z)\right]\phi_{1,yz}+{}\\ \overline{\varPhi_{11}}(Y,Z)\,\phi_{2,z}.\end{split}$ (2.28)
The second and third line here are nothing else than the previous-order normalized solutions of the Laplace equation multiplied by $\phi_{1,yz}$ instead of $\phi_{0,yz}$, and $\phi_{2,z}$ instead of $\phi_{0,z}$. The first line, once inserted into the Laplace equation for $\varPhi_{3}$, gives
$\Delta_{2}\overline{\varPhi_{31}}+2\overline{\varPhi_{21}}_{,Y}+\overline{\varPhi_{11}}=0$ (2.29)
as the determining equation for the new periodic function $\overline{\varPhi_{31}}$, with boundary conditions
$\overline{\varPhi_{31}}[Y,F(Y)]=0,\quad\lim_{Z\rightarrow\infty}\left[\overline{\varPhi_{31}}_{,Z}+Z^{2}/2+h_{1}Z\right]=0.$
From (2.29) and the already known asymptotic behaviour of $\overline{\varPhi_{21}}_{,Y}$ and $\overline{\varPhi_{11}}$ one can get the asymptotic behaviour for $Z\rightarrow\infty$ of $\overline{\varPhi_{31}}$ as
$\overline{\varPhi_{31}}\sim-Z^{3}/6-h_{1}Z^{2}/2+h_{3}.$
which, in addition to making (2.27) satisfied, of course implies the definition
$h_{3}=\lim_{Z\rightarrow\infty}\left[\overline{\varPhi_{31}}+Z^{3}/6+h_{1}Z^{2}/2\right].$ (2.30)
Extracting $\phi_{3}(y_{0},0)$ from (2.27) then gives the sought-after third-order outer boundary condition:
$\phi_{3}(y_{0},0)=h_{1}\phi_{2,z}(y_{0},0)+h_{2}\phi_{1,yz}(y_{0},0)+h_{3}\phi_{0,yyz}(y_{0},0).$ (2.31)
2.2.5 Compound equivalent boundary condition
The above process can be continued to higher and higher orders, and for each $n$ will provide an outer boundary condition of the form
$\phi_{n}(y,0)=\sum_{i=1}^{n}h_{i}\frac{\partial^{i}\phi_{n-i}}{\partial y^{i-1}\partial z}(y,0).$ (2.32)
The boundary conditions of different orders can be multiplied by $\varepsilon^{n}$ and summed together, if desired, to get the total outer $\phi$ at the virtual wall:
$\phi(y,0)=\sum_{n=0}^{N}\varepsilon^{n}\phi_{n}(y,0)+\text{O}(\varepsilon^{N+1})=\sum_{n=1}^{N}\sum_{i=1}^{n}\varepsilon^{n}h_{i}\frac{\partial^{i}\phi_{n-i}}{\partial y^{i-1}\partial z}(y,0)+\text{O}(\varepsilon^{N+1}).$ (2.33)
This equation looks like the product of two power series. In fact, by reordering summations it can be rewritten as
$\phi(y,0)=\sum_{i=1}^{N}h_{i}\frac{\partial^{i}}{\partial y^{i-1}\partial z}\sum_{n=i}^{N}\varepsilon^{n}\phi_{n-i}(y,0)+\text{O}(\varepsilon^{N+1}).$
But $\sum_{n=i}^{N}\varepsilon^{n}\phi_{n-i}=\varepsilon^{i}\sum_{n=0}^{N-i}\varepsilon^{n}\phi_{n}=\varepsilon^{i}\phi+\text{O}(\varepsilon^{N+1})$, and therefore, within an error $\text{O}(\varepsilon^{N+1})$,
$\phi(y,0)=\sum_{i=1}^{N}\varepsilon^{i}h_{i}\frac{\partial^{i}\phi}{\partial y^{i-1}\partial z}+\text{O}(\varepsilon^{N+1}),$ (2.34)
or, in dimensional form,
$\phi^{*}(y^{*},0)=\sum_{i=1}^{N}s^{*i}h_{i}\frac{\partial^{i}\phi^{*}}{\partial y^{*i-1}\partial z^{*}}+\text{O}(\varepsilon^{N+1}).$ (2.35)
Just as (2.14) was at first order, this single implicit $N^{th}$-order boundary condition is free of any reference to the outer series expansion, and may be preferred to (2.32) for computational purposes. One can also observe that if the wall shape $Z=F(Y)$ happens to be left-right symmetric, i.e. invariant under the transformation $Y\leftarrow-Y$, only even $y$-derivatives have the same symmetry. Odd $y$-derivatives are forbidden to appear in (2.32)–(2.35), and the corresponding higher-order protrusion heights must automatically come out as zero. Notice that this does not imply that the corresponding inner solutions are zero, only that they shall have the appropriate symmetry. For instance, when riblets are symmetric $\overline{\varPhi_{21}}$ of (2.22) is an odd function of $Y$, because so is $\overline{\varPhi_{11}}_{,Y}$, and its mean value, which when $Z\rightarrow\infty$ provides $h_{2}$, must be zero.
3 Longitudinal Stokes flow past riblets
In the analysis of the viscous sublayer near riblets (Luchini et al., 1991) we want to solve Stokes’ equations:
$\boldsymbol{\nabla}^{*}\boldsymbol{\cdot}\boldsymbol{u}^{*}$ $=0,$ (3.1a)
$-\boldsymbol{\nabla}^{*}p^{*}+\mu^{*}\nabla^{*2}\boldsymbol{u}^{*}$ $=\boldsymbol{0},$ (3.1b)
with boundary conditions $\boldsymbol{u}^{*}=\boldsymbol{0}$ at the solid riblet wall with riblet period $s^{*}$, and an imposed velocity gradient for $z^{*}\rightarrow\infty$. The riblet longitudinal direction is $x^{*}$, the spanwise direction is $y^{*}$ and the wall-normal direction is $z^{*}$.
For classical straight riblets described by $z^{*}/s^{*}=F(y^{*}/s^{*})$ (independent of $x^{*}$) the system of equations (3.1) can be split into longitudinal and transverse components (see, e.g., Luchini et al., 1991). The longitudinal component is governed by a Poisson equation
$\Delta^{*}_{2}u^{*}=p^{*}_{x^{*}}/\mu^{*},$ (3.2)
where the, typically negative, pressure gradient $p^{*}_{x^{*}}$ is of an order of magnitude dictated by the outer scale, and for this reason was neglected by Bechert & Bartenwerfer (1989) and by Luchini et al. (1991). When the pressure gradient is neglected, the longitudinal velocity $u^{*}$ behaves exactly like the quantity $\phi^{*}$ of §2.
At higher than first order in $\varepsilon$, the pressure gradient can no longer be neglected and, as will be seen, it provides a correction even for constant outer conditions. For the purpose of finding this correction we shall assume the pressure gradient $p^{*}_{x^{*}}/\mu^{*}$ to be proportional to $u^{*}/L^{*2}$, so that (3.2) is balanced when both sides are expressed in outer units. In other words, the outer dimensionless pressure gradient $p_{x}=L^{*2}p^{*}_{x^{*}}/\mu^{*}u_{R}^{*}$ (where $u_{R}^{*}$ is a reference velocity that will be specified later but for the time being can be arbitrary) is of order unity. Then the dimensionless form of (3.2) on the scale of the period $s^{*}$ (i.e. in inner units) is
$\Delta_{2}U=\varepsilon^{2}p_{x},$ (3.3)
where $U=u^{*}/u_{R}^{*}$ and coordinates are $Y=y^{*}/s^{*}$ and $Z=z^{*}/s^{*}$.
3.1 Wavenumber domain (constant outer conditions)
For constant $p_{x}$ and constant velocity gradient at $Z=L=1/\varepsilon$, the solution of (3.3) can be expanded in a Fourier series, just like the solution of the Laplace equation (2.1), i.e.
$U(Y,Z)=\sum_{n=-\infty}^{\infty}\hat{U}_{n}(Z)\exp(\mathrm{i}n2\pi Y).$ (3.4)
The only difference is that now $\hat{U}_{0,ZZ}$ equals $\varepsilon^{2}p_{x}$, and instead of (2.3) the mean ($n=0$) Fourier component becomes
$\hat{U}_{0}=a_{0}+b_{0}Z+\varepsilon^{2}p_{x}Z^{2}/2.$
The coefficients $a_{0}$ and $b_{0}$ in this expression could be determined, like in §2.1 from the value of the velocity gradient at large distance. However, it must be noted that while in (2.3) $\hat{\phi}^{*}_{0,z^{*}}$ is constant with $z^{*}$, and $\hat{\phi}^{*}_{0,z^{*}}(L^{*})$ directly equals $b_{0}$, in the presence of a pressure gradient the mean velocity profile becomes a quadratic, and to determine $b_{0}=\hat{U}_{0,Z}(0)$ is more complicated. If one wants to write an equivalent far-from-the-wall boundary condition transported in $Z=0$:
$a_{0}=b_{0}h_{1}+\varepsilon^{2}p_{x}h^{(p_{x})}_{2},$ (3.5)
[…]
$h^{(p_{x})}_{2}=\frac{\hat{U}_{0}(L)}{\varepsilon^{2}p_{x}}-\frac{L^{2}}{2}$
and this value is actually independent of both $p_{x}$ and $L$.
In outer coordinates, (3.5) provides an equivalent boundary condition
$u(y,0)=\varepsilon h_{1}u_{z}(y,0)+\varepsilon^{2}p_{x}h_{2}^{(p_{x})}$ (3.6)
which includes the second-order effect of the pressure gradient. Again no higher powers of $\varepsilon$ are involved, and (3.6) is exponentially accurate, when outer conditions are constant and the solution is perfectly periodic.
3.2 Physical domain (slowly varying outer conditions)
In the presence of a slowly varying outer shear $u_{0,z}(y,1)=q(y)$ and dimensionless pressure gradient $p_{x}(y,z)$, the outer problem becomes a boundary-value problem with in addition an equivalent wall boundary condition at $z=0$, which in full generality may be written by assigning $u(y,0)$ as a functional of $u_{z}(y,0)$ and $p_{x}(y,0)$. In order to obtain the equivalent boundary condition without a priori solving the outer problem, one can observe that if $u(y,0)$, $u_{z}(y,0)$ and $p_{x}$ were all simultaneously known, the outer problem could be in principle solved as an initial-value problem by Cauchy series. As long as the solutions of the boundary-value problem and the initial-value problem are the same, it is therefore sufficient that the equivalent boundary condition be consistent with this Cauchy series.
To remind what a Cauchy series is, it is in fact a Taylor series in a neighbourhood of a generic wall point $(y_{0},0)$, where $z$-derivatives higher that the first have been replaced using the differential equation differentiated zero or more times in $z$:
$u(y,z)=u(y_{0},z)+\frac{\partial u}{\partial y}(y-y_{0})+\frac{\partial u}{\partial z}z+\frac{\partial^{2}u}{\partial y^{2}}\frac{(y-y_{0})^{2}}{2}+\frac{\partial^{2}u}{\partial y\partial z}(y-y_{0})z+\left(p_{x}-\frac{\partial^{2}u}{\partial y^{2}}\right)\frac{z^{2}}{2}+\cdots$ (3.7)
We, in fact, already did a similar replacement when, in (2.16), we obtained the second derivative $\phi_{0,zz}$ from the Laplace equation as $\phi_{0,zz}=-\phi_{0,yy}$ and thus determined it was zero.
This device was originally invented by Cauchy in order to express the general solution of an initial-value problem in time, but here it will be used to reformulate our boundary-value problem as an initial-value problem in $z$. Whereas for an elliptic equation the initial-value problem is notoriously ill-posed, the trick here will be to formally apply the Cauchy series to boundary values obtained from the two-sided boundary-value problem itself. This will allow us to reformulate the inner problem as an equivalent boundary condition, without (or before) actually solving the outer problem.
For the purposes of matched asymptotic expansions, the Cauchy series will replace the outer solution that the inner solution must match. If in (3.7) $y-y_{0}$ is replaced by $\varepsilon Y$ and $z$ by $\varepsilon Z$, (3.7) becomes
$u(y,z)=u(y_{0},0)+\varepsilon\left(\frac{\partial u}{\partial y}Y+\frac{\partial u}{\partial z}Z\right)+\varepsilon^{2}\left[\frac{\partial^{2}u}{\partial y^{2}}\frac{Y^{2}}{2}+\frac{\partial^{2}u}{\partial yz}YZ+\left(p_{x}-\frac{\partial^{2}u}{\partial y^{2}}\right)\frac{Z^{2}}{2}\right]+\cdots$ (3.8)
and can be reordered as a power series in $\varepsilon$. This is the asymptotic behaviour that the inner solution must match, term by term in $\varepsilon$ and uniformly in $Y$ (or at least for large $Y$, but then automatically for any $Y$), when $Z\rightarrow\infty$.
Since the problem depends linearly on $p_{x}$, and the equivalent boundary condition can be written as a superposition of a $u(y_{0},0)$ function of $u_{z}(y_{0},0)$ and a $u(y_{0},0)$ function of $p_{x}(y_{0},0)$, it is evident that the $u_{z}(y,0)$-dependent part will be identical to the one calculated in §2 and given by (2.25). $p_{x}(y_{0},0)$, unsurprisingly, first appears in the expansion at order $\varepsilon^{2}$ in (3.8) and multiplying $Z^{2}/2$. Therefore the corresponding inner problem is
$\Delta_{2}\overline{U_{22}}=1$ (3.9)
with boundary conditions $\overline{U_{22}}[Y,F(Y)]=0$ and $\lim_{Z\rightarrow\infty}[\overline{U_{22}}_{,Z}(Y,Z)-Z]=0$. From the solution of this inner problem one extracts
$h_{2}^{(p_{x})}=\lim_{Z\rightarrow\infty}[\overline{U_{22}}(Y,Z)-Z^{2}/2]$ (3.10)
and (2.24) acquires an additional term $h_{2}^{(p_{x})}p_{0,x}(y,0)$, consistently with (3.6). (2.25) then becomes
$u(y,0)=\varepsilon h_{1}u_{z}+\varepsilon^{2}h_{2}u_{yz}+\varepsilon^{2}h_{2}^{(p_{x})}p_{x}.$ (3.11)
Consideration of the next order will be delayed until §6 since it involves coupling the longitudinal flow with the crossflow.
4 Transverse 2D Stokes problem
In the ($y^{*}$-$z^{*}$) plane orthogonal to the riblet direction $x^{*}$, the Stokes equations reduce to
$\frac{\partial v^{*}}{\partial y^{*}}+\frac{\partial w^{*}}{\partial z^{*}}$ $=0,$ (4.1a)
$-\frac{\partial p^{*}}{\partial y^{*}}+\mu^{*}\left(\frac{\partial^{2}v^{*}}{\partial y^{*2}}+\frac{\partial^{2}v^{*}}{\partial z^{*2}}\right)$ $=0,$ (4.1b)
$-\frac{\partial p^{*}}{\partial z^{*}}+\mu^{*}\left(\frac{\partial^{2}w^{*}}{\partial y^{*2}}+\frac{\partial^{2}w^{*}}{\partial z^{*2}}\right)$ $=0,$ (4.1c)
with $v^{*}=w^{*}=0$ at the riblet wall. By use of the streamfunction $\psi^{*}$ such that $v^{*}=\partial\psi^{*}/\partial z^{*}$, $w^{*}=-\partial\psi^{*}/\partial y^{*}$, continuity is automatically satisfied, and system (4.1) reduces to the biharmonic equation
$\left(\frac{\partial^{2}}{\partial y^{*2}}+\frac{\partial^{2}}{\partial z^{*2}}\right)^{2}\psi^{*}=0.$ (4.2)
This is a $4^{th}$-order elliptic equation requiring 4 boundary conditions, 2 at each side of a strip-like domain in $z^{*}$. At the solid riblet wall we have $\partial\psi^{*}/\partial z^{*}=\partial\psi^{*}/\partial y^{*}=0$, which is the same as $\partial\psi^{*}/\partial n^{*}=0$ (normal to the riblet wall, i.e. no slip) and $\partial\psi^{*}/\partial s^{*}=0$ (parallel to the riblet wall, i.e. impermeable). The latter can be integrated to $\psi^{*}=\psi^{*}_{\text{w}}$, an arbitrary constant which on one of the two walls can without loss of generality be taken as $\psi^{*}_{\text{w}}=0$.
Aim is to get 2 equivalent/homogenised boundary conditions at a reference plane $z^{*}=0$, say at the riblet crest, the most general linear form of which (compare (2.11)) can be written in matrix form as
$\left[\begin{matrix}\psi^{*}_{z^{*}}(y^{*},0)\\ \psi^{*}(y^{*},0)\end{matrix}\right]=\left[\begin{matrix}a^{*}&b^{*}\\ c^{*}&d^{*}\\ \end{matrix}\right]\left[\begin{matrix}\psi^{*}_{z^{*}z^{*}}(y^{*},0)\\ \psi^{*}_{z^{*}z^{*}z^{*}}(y^{*},0)\end{matrix}\right].$ (4.3)
These are written so as to provide two generic linear constraints among $\psi^{*}$ and its first three normal derivatives at the wall; making these constraints explicit in $\psi^{*}_{z^{*}}(y^{*},0)$ and $\psi^{*}(y^{*},0)$ has the additional advantage that $a^{*}=b^{*}=c^{*}=d^{*}=0$ corresponds to the standard conditions on a flat solid wall.
A crucial observation, which marks an important difference with the scalar case, is that the $a^{*},b^{*},c^{*},d^{*}$ coefficients have different dimensions, and each behave differently upon rescaling. If dimensionless versions of these coefficients are defined in a setting where lengths are scaled with the riblet period $s^{*}$,
$\left[\begin{matrix}s^{*}\psi^{*}_{z^{*}}(y,0)\\ \psi^{*}(y^{*},0)\end{matrix}\right]=\underbrace{\left[\begin{matrix}a_{1}&b_{2}\\ c_{2}&d_{3}\\ \end{matrix}\right]}_{\text{dimensionless}}\left[\begin{matrix}s^{*2}\psi^{*}_{z^{*}z^{*}}(y^{*},0)\\ s^{*3}\psi^{*}_{z^{*}z^{*}z^{*}}(y^{*},0)\end{matrix}\right],$
and then powers of $s^{*}$ are moved inside the center block, (4.3) becomes
$\left[\begin{matrix}\psi^{*}_{z^{*}}(y^{*},0)\\ \psi^{*}(y^{*},0)\end{matrix}\right]=\left[\begin{matrix}a_{1}s^{*}&b_{2}s^{*2}\\ c_{2}s^{*2}&d_{3}s^{*3}\\ \end{matrix}\right]\left[\begin{matrix}\psi^{*}_{z^{*}z^{*}}(y^{*},0)\\ \psi^{*}_{z^{*}z^{*}z^{*}}(y^{*},0)\end{matrix}\right],$ (4.4)
where the subscripts $1,2,3$ of the dimensionless coefficients $a_{1},b_{2},c_{2},d_{3}$ have been assigned according to their scaling.
Because of their different scaling, which is clearly tied to the simultaneous presence of derivatives of different order in the boundary conditions, the $a_{1},b_{2},c_{2},d_{3}$ coefficients will turn up at different orders in the $\varepsilon$-expansion, and only the $a_{1}$ coefficient appears at first order. Since the condition $\psi^{*}_{z^{*}}=a_{1}s^{*}\psi^{*}_{z^{*}z^{*}}$ is the same as $v^{*}=a_{1}s^{*}v^{*}_{z^{*}}$, $a_{1}$ can easily be recognized to coincide with the scalar, first-order, transverse protrusion height originally introduced in Luchini et al. (1991). A complete expansion accounting for higher-order terms will be provided in the following §4.2.
4.1 Wavenumber domain (constant outer conditions)
In Fourier space, the solution to the biharmonic equation is
$\psi^{*}(y^{*},z^{*})=\sum_{n=-\infty}^{\infty}\hat{\psi}^{*}_{n}(z^{*})\exp(\mathrm{i}n\beta^{*}_{1}y^{*}),$ (4.5)
where the far-field (mean, $n=0$) component is a third-degree polynomial given in (5) of Luchini et al. (1991),
$\begin{split}&\hat{\psi}^{*}_{0}=A^{*}_{0}+B^{*}_{0}z^{*}+C^{*}_{0}z^{*2}+D^{*}_{0}z^{*3}={}\\ &\hat{\psi}^{*}_{0}(L^{*})+\hat{\psi}^{*}_{0,z^{*}}(L^{*})(z^{*}-L^{*})+\hat{\psi}^{*}_{0,z^{*}z^{*}}(L^{*})\frac{1}{2}(z^{*}-L^{*})^{2}+\hat{\psi}^{*}_{0,z^{*}z^{*}z^{*}}(L^{*})\frac{1}{6}(z^{*}-L^{*})^{3},\end{split}$ (4.6)
and all the other components decay at least as fast as $z^{*}\exp(-\beta^{*}_{1}z^{*})=z^{*}\exp(-2\pi z^{*}/s^{*})$ with $z^{*}\rightarrow\infty$.
When the problem is linear and constant-coefficient, the $a^{*},b^{*},c^{*},d^{*}$ coefficients can be exactly recovered from a set of two numerical calculations over a domain $0\leq z^{*}\leq L^{*}$ (with $L^{*}\gg s^{*}$):
  1. 1.
    Set $\psi^{*}_{z^{*}z^{*}}(y^{*},L^{*})=1$ (arbitrary constant), $\psi^{*}_{z^{*}z^{*}z^{*}}(y^{*},L^{*})=0$, $\psi^{*}_{n^{*}}|_{\text{w}}=0$ (no slip), $\psi^{*}|_{\text{w}}=0$ (impermeable), and measure $\psi^{*}(y^{*},L^{*})$, $\psi_{z}(y^{*},L^{*})$ (constants with $y^{*}$), so we know $\boldsymbol{p}_{1}^{T}|_{L}\equiv(\hat{\psi}^{*}_{0},\hat{\psi}^{*}_{0,z^{*}},\hat{\psi}^{*}_{0,z^{*}z^{*}},\hat{\psi}^{*}_{0,z^{*}z^{*}z^{*}})_{1}(L^{*})$.
  2. 2.
    Set $\psi^{*}_{z^{*}z^{*}}(y^{*},L^{*})=0$, $\psi^{*}_{z^{*}z^{*}z^{*}}(y^{*},L^{*})=1$ (arbitrary constant), $\psi^{*}_{n^{*}}|_{\text{w}}=0$ (no slip), $\psi^{*}|_{\text{w}}=0$ (impermeable), and measure $\psi^{*}(y^{*},L^{*})$, $\psi_{z}(y^{*},L^{*})$ (constants with $y^{*}$), so we know $\boldsymbol{p}_{2}^{T}|_{L}\equiv(\hat{\psi}^{*}_{0},\hat{\psi}^{*}_{0,z^{*}},\hat{\psi}^{*}_{0,z^{*}z^{*}},\hat{\psi}^{*}_{0,z^{*}z^{*}z^{*}})_{2}(L^{*})$.
We can get the far-field solution (4.6) and its derivatives propagated from $z^{*}=L^{*}$ to $z^{*}=0$ by
$\underbrace{\left[\begin{matrix}\hat{\psi}^{*}_{0}\\ \hat{\psi}^{*}_{0,z^{*}}\\ \hat{\psi}^{*}_{0,z^{*}z^{*}}\\ \hat{\psi}^{*}_{0,z^{*}z^{*}z^{*}}\end{matrix}\right](0)}_{\boldsymbol{p}|_{0}}=\left[\begin{matrix}1&-L^{*}&(-L^{*})^{2}/2&(-L^{*})^{3}/6\\ 0&1&-L^{*}&(-L^{*})^{2}/2\\ 0&0&1&-L^{*}\\ 0&0&0&1\\ \end{matrix}\right]\underbrace{\left[\begin{matrix}\hat{\psi}^{*}_{0}\\ \hat{\psi}^{*}_{0,z^{*}}\\ \hat{\psi}^{*}_{0,z^{*}z^{*}}\\ \hat{\psi}^{*}_{0,z^{*}z^{*}z^{*}}\end{matrix}\right](L^{*})}_{\boldsymbol{p}|_{L}},$ (4.7)
so we now know how to calculate $\boldsymbol{p}_{1}|_{0}$ and $\boldsymbol{p}_{2}|_{0}$, which are linearly independent. From evaluating the two boundary-condition equations in (4.3) for each one of $\boldsymbol{p}_{1}|_{0}$ and $\boldsymbol{p}_{2}|_{0}$, we can write the four equations
$\left[\begin{matrix}(\hat{\psi}^{*}_{0,z^{*}z^{*}})_{1}|_{0}&(\hat{\psi}^{*}_{0,z^{*}z^{*}z^{*}})_{1}|_{0}\\ (\hat{\psi}^{*}_{0,z^{*}z^{*}})_{2}|_{0}&(\hat{\psi}^{*}_{0,z^{*}z^{*}z^{*}})_{2}|_{0}\end{matrix}\right]\left[\begin{matrix}a^{*}\\ b^{*}\end{matrix}\right]$ $=\left[\begin{matrix}(\hat{\psi}^{*}_{0,z^{*}})_{1}|_{0}\\ (\hat{\psi}^{*}_{0,z^{*}})_{2}|_{0}\end{matrix}\right],$
$\left[\begin{matrix}(\hat{\psi}^{*}_{0,z^{*}z^{*}})_{1}|_{0}&(\hat{\psi}^{*}_{0,z^{*}z^{*}z^{*}})_{1}|_{0}\\ (\hat{\psi}^{*}_{0,z^{*}z^{*}})_{2}|_{0}&(\hat{\psi}^{*}_{0,z^{*}z^{*}z^{*}})_{2}|_{0}\end{matrix}\right]\left[\begin{matrix}c^{*}\\ d^{*}\end{matrix}\right]$ $=\left[\begin{matrix}(\hat{\psi}^{*}_{0})_{1}|_{0}\\ (\hat{\psi}^{*}_{0})_{2}|_{0}\end{matrix}\right],$
and then solve for the four unknowns $a^{*},b^{*},c^{*},d^{*}$. The resulting equivalent boundary condition (4.4), containing terms up to third order, will be exponentially accurate, and never contain powers of $\varepsilon$ higher than the third, for constant outer boundary conditions.
4.2 Physical domain (slowly varying outer conditions)
Just like in §2.2, the dimensionless dependent variable $\psi$ can be expanded in inner coordinates $Y=(y^{*}-y_{0}^{*})/s^{*}$ and $Z=z^{*}/s^{*}$, $s^{*}$ being the wall-texture period:
$\psi=\varPsi_{0}(Y,Z)+\varepsilon\varPsi_{1}(Y,Z)+\cdots=\sum_{i=0}^{\infty}\varepsilon^{i}\varPsi_{i}(Y,Z),$ (4.8)
and outer coordinates $y=y^{*}/L^{*}$ and $z=z^{*}/L^{*}$, $L^{*}$ being an outer reference length (typically, the distance to the opposing wall):
$\psi=\psi_{0}(y,z)+\varepsilon\psi_{1}(y,z)+\cdots=\sum_{i=0}^{\infty}\varepsilon^{i}\psi_{i}(y,z).$ (4.9)
The expansion parameter $\varepsilon$ is the length ratio $s^{*}/L^{*}$.
Because the biharmonic equation is linear and scale-invariant, each term of both the inner and outer expansions will independently obey its own biharmonic equation:
$\Delta_{2}\Delta_{2}\varPsi_{i}(Y,Z)=0,\quad\Delta_{2}\Delta_{2}\psi_{i}(y,z)=0\quad\forall i.$
In addition, $\varPsi_{i}$ will obey solid-wall boundary conditions on a curve $Z=F(Y)$,
$\varPsi_{i}[Y,F(Y)]=\frac{\partial\varPsi_{i}}{\partial n}[Y,F(Y)]=0,$
with $F(Y)$ periodic of period $1$ in inner coordinates, and $\psi_{i}$ will obey two generic boundary conditions at an outer wall $z=1$.
The problem will be completed by matching conditions imposed in an intermediate distinguished limit where $Z\rightarrow\infty$ and $z\rightarrow 0$ as $\varepsilon\rightarrow 0$.
4.2.1 Order zero
At outer order 0 we need to solve the biharmonic equation for $\psi_{0}(y,z)$ with whatever boundary conditions are imposed at the outer wall and $\psi_{0}(y,0)=\psi_{0,z}(y,0)=0$ at the flat inner wall $z=0$, to which the periodic wall reduces for $\varepsilon=0$. From the (probably numerical, when the outer boundary conditions are varying in a generic way) solution of this equation we obtain, among other things, the values of $\psi_{0,zz}(y,0)$ and $\psi_{0,zzz}(y,0)$.
With $\psi(y,0)$, $\psi_{z}(y,0)$, $\psi_{zz}(y,0)$, and $\psi_{zzz}(y,0)$ known, we can formally reconstruct the outer solution by Cauchy series (actually by Taylor series until, at fourth order, we begin to see the effect of the biharmonic differential equation in eliminating fourth and higher $z$-derivatives):
$\begin{split}&\psi(y,z)=\psi(y_{0},0)+(y-y_{0})\psi_{y}(y_{0},0)+z\psi_{z}(y_{0},0)+{}\\ &\frac{(y-y_{0})^{2}}{2}\psi_{yy}(y_{0},0)+(y-y_{0})z\psi_{yz}(y_{0},0)+\frac{z^{2}}{2}\psi_{zz}(y_{0},0)+{}\\ &\frac{(y-y_{0})^{3}}{6}\psi_{yyy}(y_{0},0)+\frac{(y-y_{0})^{2}z}{2}\psi_{yyz}(y_{0},0)+\frac{(y-y_{0})z^{2}}{2}\psi_{yzz}(y_{0},0)+\frac{z^{3}}{6}\psi_{zzz}(y_{0},0)+{}\\ &\frac{z^{4}}{24}\left[-\psi_{yyyy}(y_{0},0)-2\psi_{yyzz}(y_{0},0)\right]+\cdots\end{split}$ (4.10)
For the purpose of matching, we insert the outer expansion (4.9) into (4.10), transform it in inner variables $Y=(y-y_{0})/\varepsilon$, $Z=z/\varepsilon$, and reorder in powers of $\varepsilon$:
[…]
The cancellations are due to $\psi_{0}$ and $\psi_{0,z}$ being zero, together with their $y$-derivatives, on the flat wall that corresponds to $\varepsilon=0$. This is the sequence that must be matched, term by term according to powers of $\varepsilon$, to the inner expansion (4.8).
The general matching principle is that the two higher $Z$-derivatives, $\partial^{2}\varPsi_{i}/\partial Z^{2}$ and $\partial^{3}\varPsi_{i}/\partial Z^{3}$, associated with the dominant terms in the asymptotic expansion at large $Z$, are set by the outer solution and drive the inner solution as boundary conditions in the limit of $Z\rightarrow\infty$, just as $\partial\varPhi_{i}/\partial Z$ did in §2.2, whereas the two lower $z$-derivatives, $\psi_{i}$ and $\partial\psi_{i}/\partial z$, are set by the inner solution and drive the outer solution at $z=0$, just as $\phi_{i}$ did in §2.2.
At inner order 0, matching yields all zero boundary conditions on $\varPsi_{0}(Y,Z)$; as a result $\varPsi_{0}$ is identically null (and the inner expansion could just as well have started from $i=1$).
4.2.2 First order
At inner order $1$ we obtain from (4.11)
$\varPsi_{1}(y_{0};Y,Z)\sim\psi_{1}(y_{0},0).$ (4.12)
where, just as in §2.2, $y_{0}$ is to be treated as a parameter insofar as the biharmonic equation for $\varPsi_{1}$ is concerned, and symbol $\sim$ is precisely defined to mean that the difference between its l.h.s. and r.h.s. must tend to zero when $Z\rightarrow\infty$. Second and third $Z$-derivatives of $\varPsi_{1}$ are absent (i.e. zero) in the r.h.s. of (4.12), and therefore $\varPsi_{1}$, for the same reasons as $\varPsi_{0}$, is identically null.
At outer order 1, (4.12) consequently gives $\psi_{1}(y_{0},0)=0$. The biharmonic equation, however, requires two boundary conditions at each end, and in order to find the second boundary condition for $\psi_{1}$ we need to simultaneously examine the $\varepsilon^{2}$ term of (4.11):
$\varPsi_{2}(y_{0};Y,Z)\sim\frac{Z^{2}}{2}\psi_{0,zz}+Z\psi_{1,z}+\psi_{2}(y_{0},0)$ (4.13)
where $Y\psi_{1,y}(y_{0},0)$ has been omitted because $\psi_{1}(y,0)=0$ for all $y$ and its derivative is thus also zero.
According to (4.13) the second $Z$-derivative of $\varPsi_{2}(y_{0};Y,Z)$ is non-zero at infinity, and therefore $\varPsi_{2}$ is the first non-trivial term of the inner expansion. It must be found as a solution of the biharmonic equation with the boundary conditions
$\varPsi_{2}[Y,F(Y)]=\varPsi_{2,N}[Y,F(Y)]=0,\quad\varPsi_{2,ZZ}[Y,\infty]=\psi_{0,zz}(y_{0},0),\quad\varPsi_{2,ZZZ}[Y,\infty]=0$
or equivalently, because of linearity, by means of a normalized $\overline{\varPsi_{21}}$ obeying the biharmonic equation with boundary conditions
$\overline{\varPsi_{21}}[Y,F(Y)]=\overline{\varPsi_{21}}_{,N}[Y,F(Y)]=0,\quad\overline{\varPsi_{21}}_{,ZZ}[Y,\infty]=1,\quad\overline{\varPsi_{21}}_{,ZZZ}[Y,\infty]=0,$ (4.14)
as
$\varPsi_{2}(y_{0};Y,Z)=\psi_{0,zz}(y_{0},0)\,\overline{\varPsi_{21}}(Y,Z).$ (4.15)
Once $\overline{\varPsi_{21}}$ has been determined, the first derivative of (4.13) provides the second wall boundary condition for the first-order outer solution $\psi_{1}$,
$\psi_{1,z}(y_{0},0)=\lim_{Z\rightarrow\infty}\left[\varPsi_{2,Z}-Z\psi_{0,zz}(y_{0},0)\right]=a_{1}\psi_{0,zz}(y_{0},0),$ (4.16)
where
$a_{1}=\lim_{Z\rightarrow\infty}\left(\overline{\varPsi_{21}}_{,Z}-Z\right).$ (4.17)
Constant $a_{1}$, which coincides with $a_{1}$ of §4.1 and is calculated in an equivalent manner, is the transverse protrusion height of Luchini et al. (1991). The present matching method embeds it into an expansion that can be continued to higher and higher orders.
4.2.3 Second order
Matching condition (4.13) did not yet exhaust its function. Enforcing (4.13) directly, in addition to its derivative, yields one of the boundary conditions for the second-order outer solution $\psi_{2}$:
$\psi_{2}(y_{0},0)=\lim_{Z\rightarrow\infty}\left[\varPsi_{2}(y_{0};Y,Z)-\frac{Z^{2}}{2}\psi_{0,zz}-a_{1}Z\psi_{0,zz}\right]=c_{2}\psi_{0,zz}(y_{0},0),$ (4.18)
where
$c_{2}=\lim_{Z\rightarrow\infty}\left[\overline{\varPsi_{21}}(Y,Z)-\frac{Z^{2}}{2}-a_{1}Z\right].$ (4.19)
The other boundary condition for $\psi_{2}$, as should by now be expected, is obtained from the third-order matching condition
$\varPsi_{3}\sim\frac{YZ^{2}}{2}\psi_{0,yzz}+\frac{Z^{3}}{6}\psi_{0,zzz}+YZ\psi_{1,yz}+\frac{Z^{2}}{2}\psi_{1,zz}+Y\psi_{2,y}+Z\psi_{2,z}+\psi_{3}(y_{0},0).$ (4.20)
Preliminarily the third-order inner solution $\varPsi_{3}$ must be calculated, and it can be seen to be the superposition of several contributions. The first term proportional to $\psi_{0,yzz}$ can be combined with the third using (4.16) and the fifth using (4.18), to give $Y(Z^{2}/2+a_{1}Z+c_{2})\psi_{0,yzz}$. This asymptotic behaviour can be satisfied, in analogy to (2.17), by a compound solution of the form
$\varPsi_{31}=\left[\overline{\varPsi_{31}}(Y,Z)+Y\overline{\varPsi_{21}}(Y,Z)\right]\psi_{0,yzz},$
where $\overline{\varPsi_{21}}(Y,Z)$ was defined in (4.14), and $\overline{\varPsi_{31}}(Y,Z)$ is a periodic function of $Y$ obeying the equation
$\Delta_{2}\Delta_{2}\overline{\varPsi_{31}}+4\Delta_{2}\overline{\varPsi_{21}}_{,Y}=0$ (4.21)
with boundary conditions
$\overline{\varPsi_{31}}[Y,F(Y)]=\overline{\varPsi_{31}}_{,N}[Y,F(Y)]=0,\quad\overline{\varPsi_{31}}_{,ZZ}[Y,\infty]=\overline{\varPsi_{31}}_{,ZZZ}[Y,\infty]=0.$
The second term of (4.20) proportional to $Z^{3}/6$ can be matched by a normalized contribution $\overline{\varPsi_{32}}$ which obeys the biharmonic equation with boundary conditions
$\overline{\varPsi_{32}}[Y,F(Y)]=\overline{\varPsi_{32}}_{,N}[Y,F(Y)]=0,\quad\overline{\varPsi_{32}}_{,ZZ}[Y,\infty]\sim Z,\quad\overline{\varPsi_{32}}_{,ZZZ}[Y,\infty]=1.$ (4.22)
The fourth term of (4.20) is analogous to the term multiplying $\psi_{0,zz}$ in (4.13), and when used as a boundary condition produces the same normalized solution $\overline{\varPsi_{21}}$. The sixth and seventh terms do not contribute to the second or third $Z$-derivatives, and thus have no role in determining $\varPsi_{3}$.
Combining all three contributions together provides
$\varPsi_{3}=\left[\overline{\varPsi_{31}}(Y,Z)+Y\overline{\varPsi_{21}}(Y,Z)\right]\psi_{0,yzz}+\overline{\varPsi_{32}}(Y,Z)\psi_{0,zzz}+\overline{\varPsi_{21}}(Y,Z)\psi_{1,zz}$ (4.23)
Injecting $\varPsi_{3}$ back into the $Z$-derivative of (4.20) now gives
$\psi_{2,z}(y_{0},0)=a_{2}\psi_{0,yzz}+b_{2}\psi_{0,zzz}+a_{1}\psi_{1,zz},$ (4.24)
where
$a_{2}=\lim_{Z\rightarrow\infty}\overline{\varPsi_{31}}_{,Z}(Y,Z),\quad\ b_{2}=\lim_{Z\rightarrow\infty}\left(\overline{\varPsi_{32}}_{,Z}-\frac{Z^{2}}{2}\right).$ (4.25)
Equation (4.24) closes the problem for the second-order outer solution $\psi_{2}$. Constants $b_{2}$ in (4.25) and $c_{2}$ in (4.19) have been so defined they coincide with $b_{2}$ and $c_{2}$ of §4.1. $a_{2}$ on the other hand, the analog here of $h_{2}$ in §2.2, did not exist in §4.1, and can only be discovered by matched asymptotic expansions when $\psi_{0,zz}$ is a non-constant function of $y$. Owing to variational properties of the biharmonic equation, discussed in appendix A, it also turns out that $b_{2}=-c_{2}$.
4.2.4 Third order
Just as happened for the second order, enforcing (4.20) as is yields one of the boundary conditions for the third-order outer solution $\psi_{3}$:
$\begin{split}\psi_{3}(y_{0},0)=&\lim_{Z\rightarrow\infty}\left[\varPsi_{3}-\frac{YZ^{2}}{2}\psi_{0,yzz}-\frac{Z^{3}}{6}\psi_{0,zzz}-\frac{Z^{2}}{2}\psi_{1,zz}-Y\psi_{2,y}-Z\psi_{2,z}\right]=\\ &c_{3}\psi_{0,yzz}+c_{2}\psi_{1,yzz}+d_{3}\psi_{0,zzz},\end{split}$ (4.26)
where
$c_{3}=\lim_{Z\rightarrow\infty}\left[\overline{\varPsi_{31}}(Y,Z)-a_{2}Z\right],$ (4.27)
$d_{3}=\lim_{Z\rightarrow\infty}\left[\overline{\varPsi_{32}}(Y,Z)-\frac{Z^{3}}{6}-b_{2}Z\right].$ (4.28)
Constant $d_{3}$, just as well as its previous sister constants, equals $d_{3}$ of §4.1.
The other boundary condition, specifying $\psi_{3,z}(y_{0},0)$, must be obtained from the fourth-order matching condition
$\begin{split}&\varPsi_{4}(y_{0};Y,Z)\sim\left(\frac{Y^{2}Z^{2}}{4}-\frac{Z^{4}}{12}\right)\psi_{0,yyzz}+\frac{YZ^{3}}{6}\psi_{0,yzzz}+\frac{Y^{2}Z}{2}\psi_{1,yyz}+\frac{YZ^{2}}{2}\psi_{1,yzz}+{}\\ &\frac{Z^{3}}{6}\psi_{1,zzz}+\frac{Y^{2}}{2}\psi_{2,yy}+YZ\psi_{2,yz}+\frac{Z^{2}}{2}\psi_{2,zz}+Y\psi_{3,y}+Z\psi_{3,z}+\psi_{4}(y_{0},0).\end{split}$
Two terms of this expression can be simplified using the previous boundary conditions, $\psi_{1,z}=a_{1}\psi_{0,zz}$, $\psi_{2}=c_{2}\psi_{0,zz}$, to give
$\begin{split}&\varPsi_{4}(y_{0};Y,Z)\sim\left[\frac{Y^{2}}{2}\left(\frac{Z^{2}}{2}+a_{1}Z+c_{2}\right)-\frac{Z^{4}}{12}\right]\psi_{0,yyzz}+\frac{YZ^{3}}{6}\psi_{0,yzzz}+{}\\ &\frac{YZ^{2}}{2}\psi_{1,yzz}+\frac{Z^{3}}{6}\psi_{1,zzz}+YZ\psi_{2,yz}+\frac{Z^{2}}{2}\psi_{2,zz}+Y\psi_{3,y}+Z\psi_{3,z}+\psi_{4}(y_{0},0).\end{split}$ (4.29)
In addition, the second line of (4.29) can be recognized to have the same structure as $\varPsi_{3}$ of (4.20), with only a shift from $\psi_{0}$ to $\psi_{1}$, from $\psi_{1}$ to $\psi_{2}$ and from $\psi_{3}$ to $\psi_{4}$, and the corresponding terms of $\varPsi_{4}$ can be written out accordingly.
The inner solution $\varPsi_{41}$ corresponding to the first term of (4.29) has the structure
$\varPsi_{41}=\left[\overline{\varPsi_{41}}(Y,Z)+Y\overline{\varPsi_{31}}(Y,Z)+\frac{Y^{2}}{2}\overline{\varPsi_{21}}(Y,Z)\right]\psi_{0,yyzz},$
where $\overline{\varPsi_{21}}$ and $\overline{\varPsi_{31}}$ were defined earlier, and $\overline{\varPsi_{41}}(Y,Z)$ is a periodic function of $Y$ obeying the equation
$\Delta_{2}\Delta_{2}\overline{\varPsi_{41}}+4\Delta_{2}\overline{\varPsi_{31}}_{,Y}+6\overline{\varPsi_{21}}_{,YY}+2\overline{\varPsi_{21}}_{,ZZ}=0$
with boundary conditions
$\overline{\varPsi_{41}}[Y,F(Y)]=\overline{\varPsi_{41}}_{,N}[Y,F(Y)]=0,\quad\overline{\varPsi_{41}}_{,ZZ}[Y,\infty]\sim-Z^{2},\quad\overline{\varPsi_{41}}_{,ZZZ}[Y,\infty]\sim-2Z,$
the latter two coming from the $-Z^{4}/12$ behaviour at $Z\rightarrow\infty$ expected of $\overline{\varPsi_{41}}$ in (4.29).
The second term of (4.29) asymptotically proportional to $YZ^{3}/6$ in turn can be written as
$\varPsi_{42}=\left[\overline{\varPsi_{42}}(Y,Z)+Y\overline{\varPsi_{32}}(Y,Z)\right]\psi_{0,yzzz},$
where $\overline{\varPsi_{32}}$ was defined earlier, and $\overline{\varPsi_{42}}(Y,Z)$ is a periodic function of $Y$ obeying the equation
$\Delta_{2}\Delta_{2}\overline{\varPsi_{42}}+4\Delta_{2}\overline{\varPsi_{32}}_{,Y}=0$
with boundary conditions
$\overline{\varPsi_{41}}[Y,F(Y)]=\overline{\varPsi_{41}}_{,N}[Y,F(Y)]=0,\quad\overline{\varPsi_{42}}_{,ZZ}[Y,\infty]=\overline{\varPsi_{42}}_{,ZZZ}[Y,\infty]=0,$
the latter two coming from the lack of any $Y$-independent asymptotic part in the second term of (4.29).
Combining all contributions together provides
$\begin{split}\varPsi_{4}=&\left[\overline{\varPsi_{41}}(Y,Z)+Y\overline{\varPsi_{31}}(Y,Z)+\frac{Y^{2}}{2}\overline{\varPsi_{21}}(Y,Z)\right]\psi_{0,yyzz}+\left[\overline{\varPsi_{42}}(Y,Z)+Y\overline{\varPsi_{32}}(Y,Z)\right]\psi_{0,yzzz}+{}\\ &\left[\overline{\varPsi_{31}}(Y,Z)+Y\overline{\varPsi_{21}}(Y,Z)\right]\psi_{1,yzz}+\overline{\varPsi_{32}}(Y,Z)\psi_{1,zzz}+\overline{\varPsi_{21}}(Y,Z)\psi_{2,zz}\end{split}$ (4.30)
Injecting $\varPsi_{4}$ back into the $Z$-derivative of (4.29) now gives
$\psi_{3,z}(y_{0},0)=a_{3}\psi_{0,yyzz}+b_{3}\psi_{0,yzzz}+a_{2}\psi_{1,yzz}+b_{2}\psi_{1,zzz}+a_{1}\psi_{2,zz},$ (4.31)
where
$a_{3}=\lim_{Z\rightarrow\infty}\left(\overline{\varPsi_{41}}_{,Z}(Y,Z)+2Z^{3}/3\right),\quad b_{3}=\lim_{Z\rightarrow\infty}\overline{\varPsi_{42}}_{,Z}(Y,Z).$ (4.32)
Equation (4.31), together with (4.26), closes the problem for the third-order outer solution $\psi_{3}$. Owing to special symmetries of the biharmonic equation, it also turns out that $b_{3}=-c_{3}$.
4.2.5 Higher orders and equivalent boundary condition
No apparent obstacle, aside from becoming overwhelmed with calculations, exists to continuing the procedure to higher and higher orders. Each inner contribution $\varPsi_{n}$ must in turn be asymptotically matched to an expression derived from the Cauchy series of the outer solution, in the precise sense that the difference between the two must tend to zero for $Z\rightarrow\infty$. This matching implies a growing number of numerical equalities, because both sides are asymptotically polynomials of growing degree and every coefficient of these polynomials must individually match for the two polynomials to coincide.
Each $n^{th}$ matching condition will contain two new contributions, one proportional to $\partial^{n}\psi_{0}/\partial y^{n-2}\partial z^{2}$ and one proportional to $\partial^{n}\psi_{0}/\partial y^{n-3}\partial z^{3}$, in addition to a sequence of other terms which mimic the $(n-1)^{th}$ matching condition with $\psi_{i}$ replaced by $\psi_{i+1}$. The two new contributions will produce two independent normalized inner solutions $\overline{\varPsi_{n1}}(Y,Z)$ and $\overline{\varPsi_{n2}}(Y,Z)$, and eventually produce constant numerical coefficients $a_{n-1}$, $b_{n-1}$, $c_{n}$, and $d_{n}$ for the $n^{th}$ outer boundary condition
$\left[\begin{matrix}\displaystyle\frac{\partial\psi_{n}}{\partial z}(y,0)\\ \psi_{n}(y,0)\end{matrix}\right]=\left[\begin{matrix}\displaystyle\sum_{i=1}^{n}a_{i}\frac{\partial^{i+1}\psi_{n-i}}{\partial y^{i-1}\partial z^{2}}(y,0)+\sum_{i=2}^{n}b_{i}\frac{\partial^{i+1}\psi_{n-i}}{\partial y^{i-2}\partial z^{3}}(y,0)\\ \displaystyle\sum_{i=2}^{n}c_{i}\frac{\partial^{i}\psi_{n-i}}{\partial y^{i-2}\partial z^{2}}(y,0)+\sum_{i=3}^{n}d_{i}\frac{\partial^{i}\psi_{n-i}}{\partial y^{i-3}\partial z^{3}}(y,0)\\ \end{matrix}\right].$ (4.33)
In a similar manner to (2.33), (4.33) can be multiplied by $\varepsilon^{n}$ and summed over $n=[0,N]$, so as to yield $\psi=\sum_{n=0}^{N}\varepsilon^{n}\psi_{n}+\mathrm{O}(\varepsilon^{N+1})$, and then recast as the formal product of two power series which only contains the total $\psi$ on the r.h.s as well as the l.h.s:
$\left[\begin{matrix}\displaystyle\frac{\partial\psi}{\partial z}(y,0)\\ \psi(y,0)\end{matrix}\right]=\left[\begin{matrix}\displaystyle\sum_{i=1}^{N}\varepsilon^{i}a_{i}\frac{\partial^{i+1}\psi}{\partial y^{i-1}\partial z^{2}}(y,0)+\sum_{i=2}^{N}\varepsilon^{i}b_{i}\frac{\partial^{i+1}\psi}{\partial y^{i-2}\partial z^{3}}(y,0)\\ \displaystyle\sum_{i=2}^{N}\varepsilon^{i}c_{i}\frac{\partial^{i}\psi}{\partial y^{i-2}\partial z^{2}}(y,0)+\sum_{i=3}^{N}\varepsilon^{i}d_{i}\frac{\partial^{i}\psi}{\partial y^{i-3}\partial z^{3}}(y,0)\\ \end{matrix}\right]+\mathrm{O}(\varepsilon^{N+1}).$ (4.34)
The implicit boundary condition (4.34), analogous to (2.34) for the Laplace problem, may be preferred in numerical computations. In dimensional form it can immediately be written using the wall-texture period $s^{*}$ in place of $\varepsilon$, as
$\left[\begin{matrix}\displaystyle\frac{\partial\psi^{*}}{\partial z^{*}}(y^{*},0)\\ \psi^{*}(y^{*},0)\end{matrix}\right]=\left[\begin{matrix}\displaystyle\sum_{i=1}^{N}s^{*i}a_{i}\frac{\partial^{i+1}\psi^{*}}{\partial y^{*i-1}\partial z^{*2}}(y^{*},0)+\sum_{i=2}^{N}s^{*i}b_{i}\frac{\partial^{i+1}\psi^{*}}{\partial y^{*i-2}\partial z^{*3}}(y^{*},0)\\ \displaystyle\sum_{i=2}^{N}s^{*i}c_{i}\frac{\partial^{i}\psi^{*}}{\partial y^{*i-2}\partial z^{*2}}(y^{*},0)+\sum_{i=3}^{N}s^{*i}d_{i}\frac{\partial^{i}\psi^{*}}{\partial y^{*i-3}\partial z^{*3}}(y^{*},0)\\ \end{matrix}\right].$ (4.35)
This is the generalization to slowly varying outer conditions of (4.4) which was valid for constant outer conditions.
Also, as was true in §2.2.5, when the geometry is left-right symmetric (4.33)–(4.35) must enforce the same symmetry, which means that no odd $y$-derivatives of $\psi$ can appear. The $a_{i},b_{i},c_{i},d_{i}$ coefficients corresponding to odd derivatives will automatically turn out to be zero in this case.
5 Homogenisation of the 2D Navier–Stokes equations
Up to this point we have only considered linear problems, but the matched asymptotic expansions also apply to nonlinear ones with relatively little modification, as will soon be seen. The wavenumber domain that was adopted in the previous sections when outer conditions were constant, on the other hand, loses applicability. In fact, in the original application of the protrusion height, the aim was to consider a high-Reynolds number and indeed a turbulent flow, with riblets embedded in the near-wall region known as the viscous sublayer. There should be no doubt that the reduction of the viscous sublayer from a Navier-Stokes to a Stokes problem is not a separate step but itself an application of matched asymptotic expansions, and must be consistently performed in one and the same process.
Flow past a riblet surface can be considered as having three characteristic lengths: the riblet period $s^{*}$, the distance to the opposing wall $L^{*}$ and a dynamic length which will be taken to be the viscous length $\ell^{*}=\nu^{*}/u^{*}_{\tau}$. For definiteness here $u^{*}_{\tau}$ will be the characteristic perturbation velocity that in a turbulent flow is usually defined from the mean wall-shear stress, but for the purpose of what follows any other reference velocity $u^{*}_{R}$ would do as well, as long as it correctly represents the order of magnitude of velocities, the essential underlying assumption being that all three components of velocity scale proportionally to each other (as is ordinarily the case in a near-wall turbulent flow but not, for instance, in a laminar boundary layer). The ratio $L^{*}/\ell^{*}=L^{*}u^{*}_{\tau}/\nu^{*}$ is one of the possible definitions of Reynolds number, and decides the type of flow. When $L^{*}\ll\ell^{*}$ the Reynolds number is low, $\ell^{*}$ becomes irrelevant, and we are in the Stokes regime dealt with in the previous sections. When $\ell^{*}\ll L^{*}$, the Reynolds number is high and we are in the Navier-Stokes, possibly turbulent, regime. Near the wall $L^{*}$ now becomes irrelevant, and the asymptotic-expansion parameter $\varepsilon$ must be chosen to be the ratio $\varepsilon=s^{*}/\ell^{*}$, the same parameter that is frequently denoted as $s^{+}$ (or riblet period in wall units).
In a schematization where turbulence is presumed to be dominated by quasi-streamwise vortices ($\partial/\partial x^{*}=0$), we can start from an assumption where the flow is governed by the unsteady 2D Navier-Stokes equations in primitive variables $v^{*}(y^{*},z^{*},t^{*})$, $w^{*}(y^{*},z^{*},t^{*})$, $p^{*}(y^{*},z^{*},t^{*})$:
$\begin{split}v^{*}_{t^{*}}+v^{*}v^{*}_{y^{*}}+w^{*}v^{*}_{z^{*}}+p^{*}_{y^{*}}/\rho^{*}&=\nu^{*}(v^{*}_{y^{*}y^{*}}+v^{*}_{z^{*}z^{*}}),\\ w^{*}_{t^{*}}+v^{*}w^{*}_{y^{*}}+w^{*}w^{*}_{z^{*}}+p^{*}_{z^{*}}/\rho^{*}&=\nu^{*}(w^{*}_{y^{*}y^{*}}+w^{*}_{z^{*}z^{*}}),\\ v^{*}_{y^{*}}+w^{*}_{z^{*}}&=0.\end{split}$ (5.1)
It is in fact reasonable that, especially near riblets where the largest variations occur in the $y^{*}$ and $z^{*}$ directions, $x^{*}$-derivatives should be of a higher order. Considerations of how high this order is, and the consequent effects of three-dimensionality, will be deferred to the next section.
In a nondimensionalization with reference length $\ell^{*}$, reference velocity $u^{*}_{\tau}$, reference time $\ell^{*}/u^{*}_{\tau}$, and reference pressure $\mu^{*}u_{\tau}^{*}/\ell^{*}=\rho^{*}u_{\tau}^{*2}$, all constant dimensional coefficients in these equations disappear and they become
$v_{t}+vv_{y}+wv_{z}+p_{y}=v_{yy}+v_{zz},\quad w_{t}+vw_{y}+ww_{z}+p_{z}=w_{yy}+w_{zz},\quad v_{y}+w_{z}=0.$ (5.2)
This system can be further reduced to a single equation, by introducing the streamfunction $\psi(y,z,t)$ such that $v=\psi_{z}$, $w=-\psi_{y}$. Continuity is then automatically satisfied, and the only non-zero component of the curl of the momentum equations gives
$\Delta_{2}\Delta_{2}\psi=\Delta_{2}\psi_{t}+\psi_{z}\Delta_{2}\psi_{y}-\psi_{y}\Delta_{2}\psi_{z}$ (5.3)
The highest derivatives in this equation are the same as in the transverse Stokes equation (4.2), and the required boundary conditions are also the same. In particular, 2 boundary conditions are required at each of the two walls of a strip-like domain, one of which is textured, and a solid wall is characterized by $\psi=\psi_{\text{w}}={}$constant (where the constant may, on one of the two walls, be assumed to be zero) and $\partial\psi/\partial n=0$.
5.1 Outer expansion
Just as in §4.2, the outer solution will see equivalent boundary conditions assigned at a virtual flat wall which is chosen as the origin $z=0$. The outer expansion is written as
$\psi(y,z,t)=\sum_{i=0}^{\infty}\varepsilon^{i}\psi_{i}(y,z,t)$ (5.4)
where the small parameter is, as already discussed, $\varepsilon=s^{*}/\ell^{*}=s^{+}$. Upon inserting (5.4) into (5.3) and collecting equal powers of $\varepsilon$, $\psi_{0}$ still obeys the same equation (5.3), but with no-slip and impermeability conditions on the virtual flat wall, $\psi_{0}(y,0,t)=\psi_{0,z}(y,0,t)=0$. From the (numerical, say) solution of the $\psi_{0}$-equation we can obtain, among other things, the values of $\psi_{0,zz}(y,0,t)$ and $\psi_{0,zzz}(y,0,t)$. With this information available, we can formally reconstruct the solution by Cauchy series. The beginning of this series will be identical to (4.10) up to third order, and only start to differ from the fourth-order term which is derived from the governing differential equation:
$\begin{split}&\psi(y,z,t)=\psi(y_{0},0)+(y-y_{0})\psi_{y}(y_{0},0)+z\psi_{z}(y_{0},0)+{}\\ &\frac{(y-y_{0})^{2}}{2}\psi_{yy}(y_{0},0)+(y-y_{0})z\psi_{yz}(y_{0},0)+\frac{z^{2}}{2}\psi_{zz}(y_{0},0)+{}\\ &\frac{(y-y_{0})^{3}}{6}\psi_{yyy}(y_{0},0)+\frac{(y-y_{0})^{2}z}{2}\psi_{yyz}(y_{0},0)+\frac{(y-y_{0})z^{2}}{2}\psi_{yzz}(y_{0},0)+\frac{z^{3}}{6}\psi_{zzz}(y_{0},0)+{}\\ &\frac{z^{4}}{24}\left[-\psi_{yyyy}(y_{0},0)-2\psi_{yyzz}(y_{0},0)+\psi_{z}\Delta_{2}\psi_{y}-\psi_{y}\Delta_{2}\psi_{z}+\Delta_{2}\psi_{t}\right]+\cdots\end{split}$ (5.5)
5.2 Inner expansion
The inner expansion is $\psi(y_{0}+\varepsilon Y,\varepsilon Z,t)=\varPsi(y_{0},t;Y,Z)$, where
$\varPsi(y_{0},t;Y,Z)=\sum_{i=0}^{\infty}\varepsilon^{i}\varPsi_{i}(y_{0},t;Y,Z).$ (5.6)
There is no rescaling in the time coordinate insofar as the riblet shape does not depend on time; in the inner solution time plays the role of a parameter just like $y_{0}$ does.
Since (5.3) is homogeneous with respect to space coordinates, when this equation is transformed in inner variables it remains visibly unchanged save the time derivative:
$\Delta_{2}\Delta_{2}\varPsi=\varepsilon^{2}\Delta_{2}\varPsi_{t}+\varPsi_{Z}\Delta_{2}\varPsi_{Y}-\varPsi_{Y}\Delta_{2}\varPsi_{Z}.$ (5.7)
(5.7) may then look like $\varPsi_{0}$ shall obey a nonlinear equation. However, it must be remembered that in §4.2 $\varPsi_{0}$ and $\varPsi_{1}$ turned out to be identically zero, and the expansion effectively started with $\varPsi_{2}$. There is a simple qualitative explanation for this behaviour: owing to the boundary conditions $\psi=\psi_{z}=0$, the streamfunction $\psi$ is O$(z^{2})$ near the wall $z=0$, and thus proportional to $\varepsilon^{2}Z^{2}$ in inner coordinates, and of the second order in $\varepsilon$. This estimate is unchanged in the Navier-Stokes problem. With the expansion (5.6) starting to be non-zero from $i=2$, (5.7) produces the following hierarchy:
$\Delta_{2}\Delta_{2}\varPsi_{2}$ $=0$ (5.8a)
$\Delta_{2}\Delta_{2}\varPsi_{3}$ $=0$ (5.8b)
$\Delta_{2}\Delta_{2}\varPsi_{4}$ $=\Delta_{2}\varPsi_{2,t}+\varPsi_{2,Z}\Delta_{2}\varPsi_{2,Y}-\varPsi_{2,Y}\Delta_{2}\varPsi_{2,Z}$ (5.8c)
[…]
It can be observed that each equation in (5.8) has the same, linear, l.h.s., nonlinearity being confined to the r.h.s. which at every stage is a known function of previous terms.
5.3 Orders 0, 1, 2
It begins to appear that, (5.5) and (5.8) being identical to the corresponding equations in §4.2 up to third order, nothing changes in the formulation up to inner order 3 and outer order 2 (we recall that outer order 3 partially depends upon inner order 4). This has a number of consequences:
  1. 1.
    (expected) All the previous literature that used the reduced, steady Stokes equations to describe the viscous sublayer of an otherwise general and unsteady flow was rational in doing so: this reduction can be justified by a formal asymptotic expansion, and indeed correctly represents the Navier–Stokes equations at leading order in feature size $\varepsilon=s^{+}$.
  2. 2.
    (maybe less expected) The approximation is not only good at leading order (inner order 2, outer order 1), but also at inner-order 3, outer order 2. Therefore, second-order outer boundary conditions (4.18) and (4.24) are linear and Stokes-like as well.
  3. 3.
    (however) The above second-order boundary condition involves several different contributions on equal footing, and may be more complex than previously thought (and even more complex it will become in §6).
The derivation of the inner solutions $\varPsi_{2}=\overline{\varPsi_{21}}\,\psi_{0,zz}$ and $\varPsi_{3}=(\overline{\varPsi_{31}}+Y\overline{\varPsi_{21}})\,\psi_{0,yzz}+\overline{\varPsi_{32}}\,\psi_{0,zzz}$ is identical to (4.14), (4.21), (4.22) and won’t be repeated. The corresponding equivalent boundary conditions for the outer problem are
First order: $\left\{\begin{array}[]{l}\psi_{z}(y,0,t)=\varepsilon a_{1}\psi_{zz}\\ \psi(y,0,t)=0\end{array}\right.$ (5.9c)
Second order: $\left\{\begin{array}[]{l}\psi_{z}(y,0,t)=\varepsilon a_{1}\psi_{zz}+\varepsilon^{2}a_{2}\psi_{yzz}+\varepsilon^{2}b_{2}\psi_{zzz}\\ \psi(y,0,t)=\varepsilon^{2}c_{2}\psi_{zz}\end{array}\right.$ (5.9f)
where numerical constants $a_{1}$, $a_{2}$, $b_{2}$ and $c_{2}=-b_{2}$ are defined in (4.17), (4.25), (4.19) respectively, and $a_{2}$ is expected to be zero for left-right-symmetric riblets as will be shown later to be the case.
It can be practically useful to reformulate (5.9) in primitive variables. Using $v=\psi_{z}$ and $w=-\psi_{y}$, a straightforward transformation of (5.9) gives
First order: $\left\{\begin{array}[]{l}v(y,0,t)=\varepsilon a_{1}v_{z}\\ w(y,0,t)=0\end{array}\right.$
Second order: $\left\{\begin{array}[]{l}v(y,0,t)=\varepsilon a_{1}v_{z}-\varepsilon^{2}a_{2}w_{zz}+\varepsilon^{2}b_{2}v_{zz}\\ w(y,0,t)=\varepsilon^{2}c_{2}w_{zz}\end{array}\right.$
This is not the most useful form, however. Whereas in (5.9) up to the third $z$-derivative of $\psi$ is allowed to appear, in primitive variables $v_{zz}$ and $w_{z}$ can be substituted from the momentum and continuity equations (5.2), and actually must be substituted in order to maintain compatibility with the Cauchy expansion in the developments to be presented later in the 3D extensions of this argument. (When the continuity equation, for instance, becomes 3-dimensional $u_{x}+v_{y}+w_{z}=0$, $v_{y}$ and $-w_{z}$ become no longer equivalent; $v_{y}$, and not $w_{z}$, is the quantity that appears on par with $u_{x}$ in the Cauchy expansion.)
In this minimal-$z$-derivative form, the equivalent boundary conditions become
First order: $\left\{\begin{array}[]{l}v(y,0,t)=\varepsilon a_{1}v_{z}\\ w(y,0,t)=0\end{array}\right.$ (5.10c)
Second order: $\left\{\begin{array}[]{l}v(y,0,t)=\varepsilon a_{1}v_{z}+\varepsilon^{2}a_{2}v_{yz}+\varepsilon^{2}b_{2}p_{y}\\ w(y,0,t)=-\varepsilon^{2}c_{2}v_{yz}\end{array}\right.$ (5.10f)
where $v_{0,zz}$ has been eliminated through the momentum equation $v_{0,zz}+v_{0,yy}=p_{0,y}$ combined with $v_{0,yy}=0$, and $w_{0,zz}$ has been eliminated from the $z$-derivative of the continuity equation, $w_{0,zz}=-v_{0,yz}$.
The first-order condition contains, as already stressed, the classical protrusion height for the wall-parallel velocity and zero for the wall-normal velocity. The wall-parallel second-order condition contains the wall-parallel pressure gradient in a similar manner to the longitudinal condition (3.11); however, just as (3.11) contains an additional second-order term proportional to $u_{yz}$, both of (5.10f) contain an additional term proportional to $v_{yz}$. One of the coefficients of these terms, $a_{2}$, vanishes for left-right-symmetric riblets (and in reality for any kind of riblets, as will be discovered later), but not $-c_{2}$ which always equals $b_{2}$ (see appendix A).
The same boundary conditions can be cast in dimensional form by simply replacing $\varepsilon$ by the riblet period $s^{*}$.
5.4 Third order
We recall from §4.2.4 that one of the third-order conditions, the one for $\psi_{3}(y,0)$, was obtained from the third-order inner solution. This condition remains unchanged in the Navier-Stokes problem, and is still
$\psi_{3}(y,0,t)=c_{2}\psi_{1,yzz}+c_{3}\psi_{0,yzz}+d_{3}\psi_{0,zzz},$ (5.11)
with constants $c_{2}$, $c_{3}$ and $d_{3}$ defined by (4.19), (4.27) and (4.28) respectively.
The other boundary condition, specifying $\psi_{3,z}(y,0,t)$, must be obtained from the fourth-order inner solution, which now obeys (5.8c), i.e., on account of (4.15):
$\begin{split}&\Delta_{2}\Delta_{2}\varPsi_{4}(y_{0},t;Y,Z)=\Delta_{2}\overline{\varPsi_{21}}(Y,Z)\,\psi_{0,zzt}(y_{0},0,t)+{}\\ &\left[\overline{\varPsi_{21}}_{,Z}(Y,Z)\Delta_{2}\overline{\varPsi_{21}}_{,Y}(Y,Z)-\overline{\varPsi_{21}}_{,Y}(Y,Z)\Delta_{2}\overline{\varPsi_{21}}_{,Z}(Y,Z)\right]\psi^{2}_{0,zz}(y_{0},0,t).\end{split}$
Since the l.h.s. of this equation is still linear, its solution can be obtained by simply adding to the expression (4.30) of $\varPsi_{4}$ the two additional contributions $\overline{\varPsi_{43}}(Y,Z)\,\psi_{0,zzt}(y_{0},0,t)+\overline{\varPsi_{44}}(Y,Z)\,\psi^{2}_{0,zz}(y_{0},0,t)$, where $\overline{\varPsi_{43}}$ obeys the equation
$\Delta_{2}\Delta_{2}\overline{\varPsi_{43}}=\Delta_{2}\overline{\varPsi_{21}}$ (5.12)
with boundary conditions
$\overline{\varPsi_{43}}[Y,F(Y)]=\overline{\varPsi_{43}}_{,N}[Y,F(Y)]=0,\quad\overline{\varPsi_{43}}_{,ZZ}[Y,\infty]\sim Z^{2}/2,\quad\overline{\varPsi_{43}}_{,ZZZ}[Y,\infty]\sim Z,$
and $\overline{\varPsi_{44}}$ obeys the equation
$\Delta_{2}\Delta_{2}\overline{\varPsi_{44}}=\overline{\varPsi_{21}}_{,Z}(Y,Z)\Delta_{2}\overline{\varPsi_{21}}_{,Y}(Y,Z)-\overline{\varPsi_{21}}_{,Y}(Y,Z)\Delta_{2}\overline{\varPsi_{21}}_{,Z}(Y,Z)$ (5.13)
with boundary conditions
$\overline{\varPsi_{44}}[Y,F(Y)]=\overline{\varPsi_{44}}_{,N}[Y,F(Y)]=0,\quad\overline{\varPsi_{44}}_{,ZZ}[Y,\infty]=\overline{\varPsi_{44}}_{,ZZZ}[Y,\infty]=0.$
Injecting $\varPsi_{4}$ back into the $Z$-derivative of the matching condition (4.29) now gives, instead of (4.31),
$\psi_{3,z}(y_{0},0)=a_{3}\psi_{0,yyzz}+b_{3}\psi_{0,yzzz}+a_{2}\psi_{1,yzz}+b_{2}\psi_{1,zzz}+a_{1}\psi_{2,zz}+e_{3}^{(v_{zt})}\psi_{0,zzt}+e_{3}^{(nl)}\psi_{0,zz}^{2},$ (5.14)
where the two new numerical constants are
$e_{3}^{(v_{zt})}=\lim_{Z\rightarrow\infty}\left[\overline{\varPsi_{43}}_{,Z}(Y,Z)-Z^{3}/6\right],\quad e_{3}^{(nl)}=\lim_{Z\rightarrow\infty}\overline{\varPsi_{44}}_{,Z}(Y,Z).$ (5.15)
In primitive-variable implicit form, the nonlinear third-order boundary condition is now
$\left\{\begin{array}[]{l}v(y,0,t)=\varepsilon a_{1}v_{z}+\varepsilon^{2}(a_{2}v_{yz}+b_{2}p_{y})+\varepsilon^{3}(a_{3}v_{yyz}+b_{3}p_{yy}+e_{3}^{(v_{zt})}v_{zt}+e_{3}^{(nl)}v_{z}^{2})\\ w(y,0,t)=-\varepsilon^{2}c_{2}v_{yz}-\varepsilon^{3}(c_{3}v_{yyz}+d_{3}p_{yy})\end{array}\right.$ (5.16)
Note: $e_{3}^{(nl)}$, together with $a_{2}$, $b_{3}$ and $c_{3}$, is one of those coefficients which are forbidden by symmetry, and thus zero, for left-right-symmetric riblets, because when $\overline{\varPsi_{21}}$ is even, the r.h.s. of (5.13) is odd-symmetric with respect to $Y$ and so is $\overline{\varPsi_{44}}$. It follows that for left-right-symmetric riblets nonlinearity only intervenes at an even higher than third order in $\varepsilon$.
6 3D Navier–Stokes equations
An actual turbulent flow will be three-dimensional, even when the riblets are aligned with the streamwise direction and effectively 2D. Since the matched asymptotic expansions for the 2D Navier–Stokes problem were eventually governed by linear inhomogeneous equations, we may hope for the same to happen in the more general 3D problem, which, in dimensionless form with reference velocity $u^{*}_{\tau}$, reference length $\ell^{*}=\nu^{*}/u^{*}_{\tau}$, reference time $\ell^{*}/u^{*}_{\tau}$, and reference pressure $\rho^{*}u_{\tau}^{*2}$, is
$u_{t}+uu_{x}+vu_{y}+wu_{z}+p_{x}$ $=u_{xx}+u_{yy}+u_{zz}$ (6.1a)
$v_{t}+uv_{x}+vv_{y}+wv_{z}+p_{y}$ $=v_{xx}+v_{yy}+v_{zz}$ (6.1b)
$w_{t}+uw_{x}+vw_{y}+ww_{z}+p_{z}$ $=w_{xx}+w_{yy}+w_{zz}$ (6.1c)
$u_{x}+v_{y}+w_{z}$ $=0.$ (6.1d)
In inner coordinates $Y,Z$ such that $y=y_{0}+\varepsilon Y$, $z=\varepsilon Z$, with $x$ and $t$ unchanged, these same equations become
$u_{YY}+u_{ZZ}$ $=\varepsilon(vu_{Y}+wu_{Z})+\varepsilon^{2}(-u_{xx}+u_{t}+uu_{x}+p_{x})$ (6.2a)
$v_{YY}+v_{ZZ}$ $=\varepsilon(vv_{Y}+wv_{Z}+p_{Y})+\varepsilon^{2}(-v_{xx}+v_{t}+uv_{x})$ (6.2b)
$w_{YY}+w_{ZZ}$ $=\varepsilon(vw_{Y}+ww_{Z}+p_{Z})+\varepsilon^{2}(-w_{xx}+w_{t}+uw_{x})$ (6.2c)
$v_{Y}+w_{Z}$ $=-\varepsilon u_{x}.$ (6.2d)
Let us now expand each of $u,v,w,p$ in a series of powers of $\varepsilon$, using capital letters for the terms of the inner expansion as in $u=\sum\varepsilon^{i}U_{i}$. As suggested by the previous sections, we may expect the inner expansions of $u,v,w$ (first derivatives of $\psi$) to start at order 1 and the inner expansion of pressure (akin to a second derivative of $\psi$) to start at order 0. We thus obtain the following hierarchy.
First order:
$U_{1,YY}+U_{1,ZZ}$ $=0,$ (6.3a)
$V_{1,YY}+V_{1,ZZ}-P_{0,Y}$ $=0,$ (6.3b)
$W_{1,YY}+W_{1,ZZ}-P_{0,Z}$ $=0,$ (6.3c)
$V_{1,Y}+W_{1,Z}$ $=0;$ (6.3d)
Second order:
$U_{2,YY}+U_{2,ZZ}$ $=P_{0,x},$ (6.4a)
$V_{2,YY}+V_{2,ZZ}-P_{1,Y}$ $=0,$ (6.4b)
$W_{2,YY}+W_{2,ZZ}-P_{1,Z}$ $=0,$ (6.4c)
$V_{2,Y}+W_{2,Z}$ $=-U_{1,x};$ (6.4d)
Third order:
$U_{3,YY}+U_{3,ZZ}$ $=P_{1,x}-U_{1,xx}+U_{1,t}+V_{1}U_{1,Y}+W_{1}U_{1,Z},$ (6.5a)
$V_{3,YY}+V_{3,ZZ}-P_{2,Y}$ $=-V_{1,xx}+V_{1,t}+V_{1}V_{1,Y}+W_{1}V_{1,Z},$ (6.5b)
$W_{3,YY}+W_{3,ZZ}-P_{2,Z}$ $=-W_{1,xx}+W_{1,t}+V_{1}W_{1,Y}+W_{1}W_{1,Z},$ (6.5c)
$V_{3,Y}+W_{3,Z}$ $=-U_{2,x}.$ (6.5d)
$\boldsymbol{\cdots\cdots\cdots}$
One can observe as much: i) each order is governed by the same, linear, system of l.h.s. equations, driven in various ways by the solutions of previous orders; ii) this system of equations neatly decouples into a 2D Poisson problem for $U_{n}$ and a steady 2D Stokes problem for $V_{n},W_{n},P_{n-1}$, exactly those problems that were studied in the previous sections of this paper; iii) with respect to the previous sections there are some new forcing terms, most notably in the continuity equation; iv) since the hierarchical problem is linear we can exploit superposition, taking the already acquired results for granted and focusing on each new additional term in turn.
6.1 First order
Nothing new to say. The first-order equations are homogeneous just as they were in the previous sections, and driven by the matching conditions only. They lead to the longitudinal and transverse protrusion heights of Luchini et al. (1991) according to (2.13) and (4.17).
6.2 Second order
Here the pressure gradient makes its appearance in the $U_{2}$-equation (6.4a), just as it was taken into account in §3. However, rather than being constant with $Y,Z$, the pressure gradient is shaped by the first-order transverse Stokes solution (6.3b–d). More precisely
$P_{0}=\overline{P_{01}}(Y,Z)v_{0,z}(x,y_{0},0,t)+p_{0}(x,y_{0},0,t)$
where $\overline{P_{01}}(Y,Z)$ satisfies the Cauchy-Riemann relations
$\overline{P_{01}}_{,Y}=(\Delta_{2}\overline{\varPsi_{21}})_{Z},\quad\overline{P_{01}}_{,Z}=-(\Delta_{2}\overline{\varPsi_{21}})_{Y}$ (6.6)
with $\overline{P_{01}}(Y,\infty)=0$, so that $P_{0}(Y,\infty)=p_{0}(x,y_{0},0,t)$, and $\overline{\varPsi_{21}}$ being defined by (4.14); $v_{0,z}(x,y_{0},0,t),p_{0}(x,y_{0},0,t)$ are the corresponding components of the zeroth-order outer solution at the wall. Therefore
$U_{2}=\overline{\varPhi_{21}}(Y,Z)u_{0,yz}(x,y_{0},0,t)+\overline{U_{22}}(Y,Z)p_{0,x}(x,y_{0},0,t)+\overline{U_{23}}(Y,Z)v_{0,xz}(x,y_{0},0,t)$ (6.7)
($p_{0,x}$ being the same as the constant pressure gradient formerly considered in §3), where $\overline{\varPhi_{21}}$ was defined in (2.17), $\overline{U_{22}}$ was defined in (3.9), and $\overline{U_{23}}$ obeys
$\Delta_{2}\overline{U_{23}}=\overline{P_{01}}$ (6.8)
with boundary conditions $\overline{U_{23}}[Y,F(Y)]=0,\;\overline{U_{23}}_{,Z}[Y,\infty]=0.$
The corresponding second-order outer boundary condition (3.6) now becomes
$u_{2}(x,y,0,t)=h_{1}u_{1,z}+h_{2}u_{0,yz}+h^{(p_{x})}_{2}p_{0,x}+h_{2}^{(v_{xz})}v_{0,xz},$ (6.9)
where $h_{1},h_{2},h^{(p_{x})}_{2}$ were defined in (2.13), (2.23), (3.10), and
$h_{2}^{(v_{xz})}=\lim_{Z\rightarrow\infty}\overline{U_{23}}(Y,Z).$ (6.10)
As to the r.h.s. of (6.4d), it makes a streamfunction impossible to use to derive the velocity field, and the simplest choice is to solve these equations numerically in primitive form. Nonetheless, their solution can be written as the linear superposition of a solution with non-zero forcing from $v_{0,z}(x,y_{0},0,t)$ and zero r.h.s., which is exactly the one already dealt with in §4.2.3, and one with $v_{0,z}(x,y_{0},0,t)=0$ and non-zero r.h.s. of (6.4d), which can be represented through a normalized solution $\overline{V_{23}},\overline{W_{23}},\overline{P_{13}}$ that solves
$\overline{V_{23}}_{,YY}+\overline{V_{23}}_{,ZZ}-\overline{P_{13}}_{,Y}$ $=0,$ (6.11a)
$\overline{W_{23}}_{,YY}+\overline{W_{23}}_{,ZZ}-\overline{P_{13}}_{,Z}$ $=0,$ (6.11b)
$\overline{V_{23}}_{,Y}+\overline{W_{23}}_{,Z}$ $=-\overline{\varPhi_{11}};$ (6.11c)
with $\overline{\varPhi_{11}}$ taken from (2.12) and boundary conditions $\overline{V_{23}}[Y,F(Y)]=\overline{W_{23}}[Y,F(Y)]=0,\;\overline{V_{23}}_{,Z}(Y,\infty)=0,\;\overline{P_{13}}(Y,\infty)\sim-Z.$
Then, since $U_{1,x}(x,y_{0},t;Y,Z)=\overline{\varPhi_{11}}(Y,Z)\,u_{0,xz}(x,y_{0},0,t)$, the second-order boundary conditions (5.10) become
$\left\{\begin{array}[]{l}v(x,y,0,t)=\varepsilon a_{1}v_{z}+\varepsilon^{2}a_{2}v_{yz}+\varepsilon^{2}b_{2}p_{y}+\varepsilon^{2}f_{2}u_{xz}\\ w(x,y,0,t)=-\varepsilon^{2}c_{2}v_{yz}+\varepsilon^{2}g_{2}u_{xz},\end{array}\right.$ (6.12)
where the two newly introduced numerical constants are
$f_{2}=\lim_{Z\rightarrow\infty}\overline{V_{23}}(Y,Z),\quad\ g_{2}=\lim_{Z\rightarrow\infty}[\overline{W_{23}}(Y,Z)+Z^{2}/2+h_{1}Z].$ (6.13)
6.3 Third order
The procedure should by now be self-evident. Each stage of the $\varepsilon$-expansion entails a growing number of contributions, which superpose linearly and can be handled independently of each other. All of the terms already included in the 2D analysis are present. As compared to the 2D Navier-Stokes equations discussed in §5.4, the third order (6.5) of the 3D Navier-Stokes problem produces the following additions.
6.3.1 Longitudinal strain
Each component of the momentum equation contains a second $x$-derivative, representing viscous longitudinal strain, in a position analogous to the time derivative, as highlighted below:
$U_{3,YY}+U_{3,ZZ}$ $=P_{1,x}\,\boxed{-U_{1,xx}+U_{1,t}}+V_{1}U_{1,Y}+W_{1}U_{1,Z},$
$V_{3,YY}+V_{3,ZZ}-P_{2,Y}$ $=\boxed{-V_{1,xx}+V_{1,t}}+V_{1}V_{1,Y}+W_{1}V_{1,Z},$
$W_{3,YY}+W_{3,ZZ}-P_{2,Z}$ $=\boxed{-W_{1,xx}+W_{1,t}}+V_{1}W_{1,Y}+W_{1}W_{1,Z},$
$V_{3,Y}+W_{3,Z}$ $=-U_{2,x}.$
Indeed, the analogy is so precise that in the transverse equivalent boundary condition (5.16) it suffices to replace $v_{0,zt}$ by $v_{0,zt}-v_{0,xxz}$ (and $w_{0,zt}$ by $w_{0,zt}-w_{0,xxz}$ if they were present; but, by continuity and the no-slip conditions on $u_{0}$ and $v_{0}$, $w_{0,zt}=w_{0,xxz}=0$). In the longitudinal momentum equation, where the time derivative was not included in the analysis of §3, this contribution to $U_{3}$ can be written as $\overline{U_{33}}(Y,Z)[u_{0,zt}(x,y_{0},0,t)-u_{0,xxz}(x,y_{0},0,t)]$, where $\overline{U_{33}}$ is the solution of
$\Delta_{2}\overline{U_{33}}=\overline{\varPhi_{11}},$ (6.14)
$\overline{\varPhi_{11}}$ being defined in (2.12), with boundary conditions $\overline{U_{33}}[Y,F(Y)]=0,\;\overline{U_{33}}_{,Z}(Y,\infty)\sim Z^{2}/2+h_{1}Z.$
There follows an additional term to the equivalent boundary condition (3.11) of the form $\varepsilon^{3}h_{3}^{(u_{zt})}(u_{zt}-u_{xxz})$, where
$h_{3}^{(u_{zt})}=\lim_{Z\rightarrow\infty}[\overline{U_{33}}(Y,Z)-Z^{3}/6-h_{1}Z^{2}/2].$ (6.15)
6.3.2 Nonlinear terms
The nonlinear contributions to the third-order inner solution,
$U_{3,YY}+U_{3,ZZ}$ $=P_{1,x}-U_{1,xx}+U_{1,t}+\boxed{V_{1}U_{1,Y}+W_{1}U_{1,Z}}\,,$
$V_{3,YY}+V_{3,ZZ}-P_{2,Y}$ $=-V_{1,xx}+V_{1,t}+\boxed{V_{1}V_{1,Y}+W_{1}V_{1,Z}}\,,$
$W_{3,YY}+W_{3,ZZ}-P_{2,Z}$ $=-W_{1,xx}+W_{1,t}+\boxed{V_{1}W_{1,Y}+W_{1}W_{1,Z}}\,,$
$V_{3,Y}+W_{3,Z}$ $=-U_{2,x},$
can be subdivided into a transverse part, which is identical to the one already studied in §5.4, and a longitudinal part which represents transport of longitudinal momentum caused by the crossflow velocities. The r.h.s. for the longitudinal part can be written as $[\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}_{,Z}]v_{0,z}u_{0,z}$. Therefore the corresponding contribution to $U_{3}$ is $\overline{U_{34}}(Y,Z)v_{0,z}u_{0,z}$, where
$\Delta_{2}\overline{U_{34}}=\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}_{,Z}$ (6.16)
with boundary conditions $\overline{U_{34}}[Y,F(Y)]=0,\;\overline{U_{34}}_{,Z}(Y,\infty)=0.$
There ensues an additional term to the equivalent boundary condition (3.11) of the form $\varepsilon^{3}h^{(nl)}_{3}v_{0,z}u_{0,z}$, where
$h^{(nl)}_{3}=\lim_{Z\rightarrow\infty}\overline{U_{34}}(Y,Z).$ (6.17)
Just as its transverse counterpart, this nonlinear contribution is forbidden (the r.h.s. of (6.16) is an odd function of $Y$ and so is $\overline{U_{34}}$) for left-right-symmetric wall features.
6.3.3 Cross-coupling
Finally, (6.5) contain the same cross-coupling terms between longitudinal and transverse components that made their first appearance at the second order:
$U_{3,YY}+U_{3,ZZ}$ $=\boxed{P_{1,x}}-U_{1,xx}+U_{1,t}+V_{1}U_{1,Y}+W_{1}U_{1,Z},$
$V_{3,YY}+V_{3,ZZ}-P_{2,Y}$ $=-V_{1,xx}+V_{1,t}+V_{1}V_{1,Y}+W_{1}V_{1,Z},$
$W_{3,YY}+W_{3,ZZ}-P_{2,Z}$ $=-W_{1,xx}+W_{1,t}+V_{1}W_{1,Y}+W_{1}W_{1,Z},$
$V_{3,Y}+W_{3,Z}$ $=\boxed{-U_{2,x}}\,.$
These cross-couplings ingenerate several third-order contributions, one for every second-order term that takes the stage on the r.h.s. in turn.
$P_{1,x}$ can be obtained from $\varPsi_{3}$ of (4.23), with additional contributions from $\overline{P_{13}}$ of (6.11) and from the order-1 outer pressure gradient at the wall. Summing the $Z$-derivative of (4.23) with the term involving $\overline{V_{23}}$ of (6.11) gives
$\begin{split}V_{2}&=\left[\overline{\varPsi_{31}}_{,Z}(Y,Z)+Y\overline{\varPsi_{21}}_{,Z}(Y,Z)\right]v_{0,yz}(x,y_{0},0,t)+\overline{\varPsi_{32}}_{,Z}(Y,Z)p_{0,y}(x,y_{0},0,t)+{}\\ &\overline{V_{23}}(Y,Z)u_{0,xz}(x,y_{0},0,t)+\overline{\varPsi_{21}}_{,Z}(Y,Z)v_{1,z}(x,y_{0},0,t)\end{split}$ (6.18)
and correspondingly
$\begin{split}P_{1,x}&=\left[\overline{P_{11}}(Y,Z)+Y\overline{P_{01}}(Y,Z)\right]v_{0,xyz}(x,y_{0},0,t)+\overline{P_{12}}(Y,Z)p_{0,xy}(x,y_{0},0,t)+{}\\ &\overline{P_{13}}(Y,Z)u_{0,xxz}(x,y_{0},0,t)+\overline{P_{01}}(Y,Z)v_{1,xz}(x,y_{0},0,t)+p_{1,x}(x,y_{0},0,t).\end{split}$ (6.19)
To convert (6.18) into (6.19) it mostly suffices to apply Cauchy-Riemann relationships analogous to (6.6) term by term. This is true, in particular, for $\overline{P_{01}}$, already available from (6.6) itself, and for $\overline{P_{12}}$, obtained from a similar set of equations applied to $\overline{\varPsi_{32}}$, whereas $\overline{P_{13}}$ directly comes from (6.11). $\overline{P_{11}}$ requires special care, owing to the presence of the term proportional to $Y$ in the square brackets.
Let us write Cauchy-Riemann relationships for the entire square brackets in (6.18) and (6.19):
$\left[\overline{P_{11}}+Y\overline{P_{01}}\right]_{Y}=\Delta_{2}\left[\overline{\varPsi_{31}}+Y\overline{\varPsi_{21}}\right]_{Z},\quad\left[\overline{P_{11}}+Y\overline{P_{01}}\right]_{Z}=-\Delta_{2}\left[\overline{\varPsi_{31}}+Y\overline{\varPsi_{21}}\right]_{Y}.$
Expanding the derivatives, and taking into account that $\overline{P_{01}},\overline{\varPsi_{21}}$ already obey their own Cauchy-Riemann relationships, gives
$\overline{P_{11}}_{,Y}+\overline{P_{01}}=\Delta_{2}\overline{\varPsi_{31}}_{,Z}+2\overline{\varPsi_{21}}_{,YZ},\quad\overline{P_{11}}_{,Z}=-\Delta_{2}\overline{\varPsi_{31}}_{,Y}-3\overline{\varPsi_{21}}_{,YY}-\overline{\varPsi_{21}}_{,ZZ}.$ (6.20)
$\overline{P_{11}}$ must be obtained from (6.20), with a boundary condition $\overline{P_{11}}(Y,\infty)\sim-Z$.
The contribution to $U_{3}$ resulting from all the terms of (6.19) together, to be added to (2.28), will be
$\begin{split}&\left[\overline{U_{35}}(Y,Z)+Y\overline{U_{23}}(Y,Z)\right]v_{0,xyz}(x,y_{0},0,t)+\left[\overline{U_{36}}(Y,Z)+Y\overline{U_{22}}(Y,Z)\right]p_{0,xy}(x,y_{0},0,t)+{}\\ &\overline{U_{37}}(Y,Z)u_{0,xxz}(x,y_{0},0,t)+\overline{U_{23}}(Y,Z)v_{1,xz}(x,y_{0},0,t)+\overline{U_{22}}p_{1,x}(x,y_{0},0,t)\end{split}$
where $\overline{U_{22}}$ and $\overline{U_{23}}$ were already defined in (3.9) and (6.8) respectively, and $\overline{U_{35}}$, $\overline{U_{36}}$, $\overline{U_{37}}$ obey
$\Delta_{2}\overline{U_{35}}+2\overline{U_{23}}_{,Y}=\overline{P_{11}},\quad\Delta_{2}\overline{U_{36}}+2\overline{U_{22}}_{,Y}=\overline{P_{12}},\quad\Delta_{2}\overline{U_{37}}=\overline{P_{13}}$
with boundary conditions $\overline{U_{35}}_{,Z}(Y,\infty)\sim-Z^{2}/2$, $\overline{U_{36}}_{,Z}(Y,\infty)=0$, $\overline{U_{37}}_{,Z}(Y,\infty)=0$.
The equivalent boundary condition for $u(x,y,0,t)$, (3.11), will then acquire new contributions
$\varepsilon^{3}h_{3}^{(v_{xyz})}v_{xyz}+\varepsilon^{3}h_{3}^{(p_{xy})}p_{xy}+\varepsilon^{3}h_{3}^{(u_{xxz})}u_{xxz}$
where
$\begin{split}&h_{3}^{(v_{xyz})}=\lim_{Z\rightarrow\infty}\left[\overline{U_{35}}(Y,Z)+Z^{3}/6\right],\quad h_{3}^{(p_{xy})}=\lim_{Z\rightarrow\infty}\overline{U_{36}}(Y,Z),\\ &h_{3}^{(u_{xxz})}=\lim_{Z\rightarrow\infty}\overline{U_{37}}(Y,Z).\end{split}$ (6.21)
As far as the r.h.s. of the continuity equation (6.5d) is concerned, the $x$-derivative of the second-order longitudinal velocity, $U_{2}$ of (6.7), is
$\begin{split}U_{2,x}&=\overline{\varPhi_{11}}(Y,Z)u_{1,xz}(x,y_{0},0,t)+\left[\overline{\varPhi_{21}}(Y,Z)+Y\,\overline{\varPhi_{1}(Y,Z)}\right]u_{0,xyz}(x,y_{0},0,t)+{}\\ &\overline{U_{22}}(Y,Z)p_{0,xx}(x,y_{0},0,t)+\overline{U_{23}}(Y,Z)v_{0,xxz}(x,y_{0},0,t).\end{split}$ (6.22)
Each term in turn must be inserted into four problems similar to (6.11), one with r.h.s. $-\overline{\varPhi_{11}}(Y,Z)$ yielding the same $\overline{V_{23}},\overline{W_{23}},\overline{P_{13}}$ as (6.11), one with r.h.s. $-\overline{\varPhi_{21}}(Y,Z)-Y\overline{\varPhi_{11}}(Y,Z)$ yielding $\overline{V_{35}},\overline{W_{35}},\overline{P_{25}}$, one with r.h.s. $-\overline{U_{22}}(Y,Z)$ yielding $\overline{V_{36}},\overline{W_{36}},\overline{P_{26}}$ and one with r.h.s. $-\overline{U_{23}}(Y,Z)$ yielding $\overline{V_{37}},\overline{W_{37}},\overline{P_{27}}$.
Just as with (6.20), the term $-\overline{\varPhi_{21}}(Y,Z)-Y\overline{\varPhi_{11}}(Y,Z)$ requires special care for the presence of a part proportional to $Y$. Inserting this binomial into a Stokes problem like (6.11) gives
$(\overline{V_{35}}+Y\overline{V_{23}})_{YY}+(\overline{V_{35}}+Y\overline{V_{23}})_{ZZ}-(\overline{P_{25}}+Y\overline{P_{13}})_{Y}$ $=0,$
$(\overline{W_{35}}+Y\overline{W_{23}})_{YY}+(\overline{W_{35}}+Y\overline{W_{23}})_{ZZ}-(\overline{P_{25}}+Y\overline{P_{13}})_{Z}$ $=0,$
$(\overline{V_{35}}+Y\overline{V_{23}})_{Y}+(\overline{W_{35}}+Y\overline{W_{23}})_{Z}$ $=-\overline{\varPhi_{21}}-Y\overline{\varPhi_{11}},$
i.e., upon remembering that $\overline{V_{23}},\overline{W_{23}},\overline{P_{13}}$ themselves satisfy (6.11),
$\overline{V_{35}}_{,YY}+\overline{V_{35}}_{,ZZ}-\overline{P_{25}}_{,Y}$ $=\overline{P_{13}}-2\overline{V_{23}}_{,Y},$ (6.23a)
$\overline{W_{35}}_{,YY}+\overline{W_{35}}_{,ZZ}-\overline{P_{25}}_{,Z}$ $=-2\overline{W_{23}}_{,Y},$ (6.23b)
$\overline{V_{35}}_{,Y}+\overline{W_{35}}_{,Z}$ $=-\overline{\varPhi_{21}}-\overline{V_{23}}.$ (6.23c)
This is the system of equations to obtain $\overline{V_{35}},\overline{W_{35}},\overline{P_{25}}$ from.
There will result three new contributions to the $v$ and $w$ equivalent boundary conditions (5.16),
$\left\{\begin{array}[]{l}\varepsilon^{3}f_{3}^{(u_{xyz})}u_{xyz}+\varepsilon^{3}f_{3}^{(p_{xx})}p_{xx}+\varepsilon^{3}f_{3}^{(v_{xxz})}v_{xxz}\\ \varepsilon^{3}g_{3}^{(u_{xyz})}u_{xyz}+\varepsilon^{3}g_{3}^{(p_{xx})}p_{xx}+\varepsilon^{3}g_{3}^{(v_{xxz})}v_{xxz}\\ \end{array}\right.$
where
$f_{3}^{(u_{xyz})}=\lim_{Z\rightarrow\infty}[\overline{V_{35}}(Y,Z)+Z^{3}/6],\quad f_{3}^{(p_{xx})}=\lim_{Z\rightarrow\infty}\overline{V_{36}}(Y,Z),\quad f_{3}^{(v_{xxz})}=\lim_{Z\rightarrow\infty}\overline{V_{37}}(Y,Z)$ (6.24)
and
$\begin{split}&g_{3}^{(u_{xyz})}=\lim_{Z\rightarrow\infty}[\overline{W_{35}}(Y,Z)+(h_{2}+f_{2})Z],\quad g_{3}^{(p_{xx})}=\lim_{Z\rightarrow\infty}[\overline{W_{36}}(Y,Z)+Z^{3}/6+h_{2}^{(p_{x})}Z],\\ &g_{3}^{(v_{xxz})}=\lim_{Z\rightarrow\infty}[\overline{W_{37}}(Y,Z)+h_{2}^{(v_{xz})}Z].\end{split}$ (6.25)
7 Summing-up and numerical examples
Collecting the results of the last five sections together, the complete, up to third-order, equivalent boundary condition obtained through matched asymptotic expansions is
$\begin{split}&\left[\begin{matrix}u\\ v\\ w\end{matrix}\right]=\ \varepsilon\left[\begin{matrix}h_{1}u_{z}\\ a_{1}v_{z}\\ 0\end{matrix}\right]+\ \varepsilon^{2}\left[\begin{matrix}{\color[rgb]{.5,.5,.5}h_{2}u_{yz}}+{}\\ {\color[rgb]{.5,.5,.5}a_{2}v_{yz}}+{}\\ {}\end{matrix}\begin{matrix}h_{2}^{(p_{x})}p_{x}+{\color[rgb]{.5,.5,.5}h_{2}^{(v_{xz})}v_{xz}}\\ b_{2}p_{y}+{\color[rgb]{.5,.5,.5}f_{2}u_{xz}}\\ -c_{2}v_{yz}+g_{2}u_{xz}\end{matrix}\right]+{}\\ &\varepsilon^{3}\left[\begin{matrix}h_{3}u_{yyz}+{\color[rgb]{.5,.5,.5}h^{(p_{xy})}_{3}p_{xy}}+h_{3}^{(u_{zt})}(u_{zt}-u_{xxz})+{\color[rgb]{.5,.5,.5}h^{(nl)}_{3}u_{z}v_{z}}+h_{3}^{(v_{xyz})}v_{xyz}+h_{3}^{(u_{xxz})}u_{xxz}\\ a_{3}v_{yyz}+{\color[rgb]{.5,.5,.5}b_{3}p_{yy}}+e_{3}^{(v_{zt})}(v_{zt}-v_{xxz})+{\color[rgb]{.5,.5,.5}e_{3}^{(nl)}v_{z}^{2}}+f_{3}^{(u_{xyz})}u_{xyz}+{\color[rgb]{.5,.5,.5}f_{3}^{(p_{xx})}p_{xx}}+f_{3}^{(v_{xxz})}v_{xxz}\\ -c_{2}v_{yz}+g_{2}u_{xz}+{\color[rgb]{.5,.5,.5}g_{3}^{(u_{xyz})}u_{xyz}}+g_{3}^{(p_{xx})}p_{xx}+{\color[rgb]{.5,.5,.5}g_{3}^{(v_{xxz})}v_{xxz}}\end{matrix}\right]\end{split}$ (7.1)
where terms that are forbidden by left-right symmetry (when the wall texture has got such a symmetry) have been greyed out.
Shape: triangular rectangular trapezoidal sawtooth slip-no-slip
$h_{1}$ (2.13) 00.170741 00.114319 00.191366 00.127310 00.110325
$a_{1}$ (4.17) 00.080200 00.048000 00.081700 00.073500 00.054700
$h_{2}$ (2.23) 00.000000 00.000000 00.000000 00.000000 00.000000
$h_{2}^{(p_{x})}$ (3.10) -0.028211 -0.021001 -0.034823 -0.014104 00.000000
$h_{2}^{(v_{xz})}$ (6.10) 00.000000 00.000000 00.000000 00.003379 00.000000
$a_{2}$ (4.25) 00.000000 00.000000 00.000000 00.000000 00.000000
$b_{2}$ (4.25) -0.005823 -0.002076 -0.005411 -0.004794 00.000000
$-c_{2}$ (4.19) -0.005823 -0.002076 -0.005411 -0.004794 00.000000
$f_{2}$ (6.13) 00.000000 00.000000 00.000000 -0.003392 00.000000
$g_{2}$ (6.13) -0.028213 -0.021001 -0.034825 -0.014105 00.000000
$h_{3}$ (2.30) 00.001658 00.000497 00.002330 00.000686 -0.000767
$h_{3}^{(u_{zt})}$ (6.15) -0.003593 -0.002356 -0.004819 -0.001505 -0.001050
$h_{3}^{(nl)}$ (6.17) 00.000000 00.000000 00.000000 00.000000 00.000000
$h_{3}^{(v_{xyz})}$ (6.21) -0.002664 -0.001329 -0.003696 -0.000959 -0.000756
$h_{3}^{(p_{xy})}$ (6.21) 00.000000 00.000000 00.000000 -0.000232 00.000000
$h_{3}^{(u_{xxz})}$ (6.21) 00.001898 00.004648 00.002032 00.000986 00.001735
$a_{3}$ (4.32) 00.003013 00.001885 00.002978 00.002348 00.002452
$b_{3}$ (4.32) 00.000000 00.000000 00.000000 -0.000199 00.000000
$-c_{3}$ (4.27) 00.000000 00.000000 00.000000 -0.000200 00.000000
$-d_{3}$ (4.28) 00.001665 00.000981 00.001975 00.000848 00.000000
$e_{3}^{(v_{zt})}$ (5.15) -0.000108 -0.000196 -0.000132 -0.000090 -0.000259
$e_{3}^{(nl)}$ (5.15) 00.000000 00.000000 00.000000 00.000000 00.000000
$f_{3}^{(u_{xyz})}$ (6.24) -0.001350 -0.000559 -0.001833 -0.000551 -0.000751
$f_{3}^{(p_{xx})}$ (6.24) 00.000000 00.000000 00.000000 00.001071 00.000000
$f_{3}^{(v_{xxz})}$ (6.24) -0.000998 -0.000769 -0.001394 -0.001029 -0.000448
$g_{3}^{(u_{xyz})}$ (6.25) 00.000000 00.000000 00.000000 00.000219 00.000000
$g_{3}^{(p_{xx})}$ (6.25) 00.010247 00.009021 00.012171 00.003065 00.000000
$g_{3}^{(v_{xxz})}$ (6.25) 00.000000 00.000000 00.000000 -0.001070 00.000000
Table 1: Up to third-order protrusion heights for some typical riblet shapes.
The coefficients involved in this formula have been computed numerically for five wall shapes, chosen for the sake of example but also for their objective interest:
  1. 1.
    Triangular riblets with an equilateral shape, tip angle $60^{\circ}$, and height $\sqrt{3/4}s^{*}$. A relatively common shape in both simulations and experiments.
  2. 2.
    Rectangular riblets with height $0.5s^{*}$ and width $0.2s^{*}$. Not particularly effective for drag reduction, but easy to mesh for the purpose of numerical simulation and therefore useful for comparison by other authors.
  3. 3.
    Trapezoidal riblets with height $0.5s^{*}$ and tip angle $30^{\circ}$, close to what is empirically considered to provide the best drag reduction with a practical shape.
  4. 4.
    Asymmetric riblets with a sawtooth shape, height $0.5s^{*}$, and tip angle such as to yield a width equal to the period $s^{*}$. Devised to test the effects of asymmetry on those coefficients of (7.1) that are mandated to be zero when riblets are left-right symmetric. Some flow in this geometry was computed by Modesti et al. (2021).
  5. 5.
    The fifth pattern is not really an indented geometry, but a pattern of slip and no-slip stripes on an impermeable flat surface, with $50\%$$50\%$ surface coverage. This configuration is the simplest idealized model of a superhydrophobic surface when used for drag reduction. It also possesses analytical solutions (Philip, 1972), which provide its two first-order protrusion heights as $h_{1}=\log(2)/(2\pi),a_{1}=h_{1}/2$.
For each of the coefficients in (7.1) a corresponding numerical viscous problem has been solved. This is a less daunting computational task than it may look like because only two solvers have to be programmed, one for a 2D Poisson problem and one for a 2D Stokes problem, with a single set of boundary conditions each, and then applied multiple times with different right-hand sides of either the equations or boundary conditions or both. The equations for either problem have been discretized on a square grid using a second-order, finite-difference, immersed-boundary method (Luchini et al., 2025), enhanced with an analytical correction for the tip singularity. The resulting discrete linear system has been LU-decomposed by direct banded Gauss elimination, just once, and then applied to a battery of different right-hand sides. The values obtained for the various coefficients are collected in table 1.
Figure 2: First-order streamwise velocity (a), spanwise velocity (b), wall-normal velocity (c), and corresponding zeroth-order pressure (d) for the five considered geometries. The riblet profile is drawn on top of each figure.
Figure 3: Second-order contributions to the streamwise velocity. (a) $\overline{\varPhi_{21}}$ of (2.18), (b) $\overline{U_{22}}$ of (3.9), (c) $\overline{U_{23}}$ of (6.8).
Figure 4: Second-order contributions to the spanwise velocity. (a) $\overline{\varPsi_{31}}_{,Z}$ of (4.21), (b) $\overline{\varPsi_{32}}_{,Z}$ of (4.22), (c) $\overline{V_{23}}$ of (6.11).
Figure 5: Third-order velocities ingenerated by the time derivative. (a) streamwise velocity $\overline{U_{33}}$ of (6.14), (b) spanwise velocity $\overline{\varPsi_{43}}_{,Z}$ of (5.12).
Figure 6: Nonlinear third-order velocities. (a) streamwise velocity $\overline{U_{34}}$ of (6.16), (b) spanwise velocity $\overline{\varPsi_{44}}_{,Z}$ of (5.13).
Figure 7: Wall-normal velocity corresponding to the largest wall-normal third-order correction, $\overline{W_{36}}$ defined just below (6.22).
A selection of corresponding velocity fields are collected in figures 27. Each plot covers a $1\times 1.5$ patch of a taller $1\times 3$ computational domain covered with $241\times 721$ grid points ($121\times 181$ displayed). The computational domain can be kept relatively short because each solution converges exponentially to its asymptotic behaviour in $Z$.
We can remark that, of the 2 first-order, 8 second-order, and 18 third-order potential coefficients, several are zero. For left-right symmetric riblets, which means for the first three columns of table 1, 2 first-order, 4 second-order and 10 third-order coefficients are allowed to be non-zero, as the rest are forbidden by symmetry (as brought out in table 1). In addition, two pairs of second-order coefficients are tied to each other ($b_{2}=-c_{2},h_{2}^{(p_{x})}=g_{2}$), thus making the number of independent second-order coefficients just 2. One can observe in figures 3 and 4 that for each forbidden coefficient the corresponding velocity field is of odd symmetry, thus making its $Y$-average zero.
But even for asymmetric riblets, in the fourth column, astoundingly enough 2 second-order coefficients and 2 third-order coefficients are zero, thus making the totals 2, 6 and 16. This is not just a coincidence or a numerical approximation: as shown in appendix B, these coefficients can actually be proved to be zero for any wall shape. They are also not just any of the terms: $h_{2}$ and $a_{2}$ are the first higher-order protrusion heights that came about in the derivation ((2.23) and (4.25) respectively), and even more importantly $e_{3}^{(nl)}$ and $h_{3}^{(nl)}$ are the only third-order terms that involve nonlinearities of the Navier-Stokes equations ((5.15) and (6.17) respectively). Therefore nonlinearities, which were already barred by symmetry for left-right symmetric wall textures, are actually absent, up to third order, for any texture.
One can also notice from the table that not all coefficients are independent, since $b_{2}=-c_{2}$, $b_{3}=-c_{3}$, $f_{2}=-h_{2}^{(v_{xz})}$, $g_{2}=h_{2}^{(p_{x})}$, $h_{3}^{(p_{xy})}=-g_{3}^{(u_{xyz})}$, and $f_{3}^{(p_{xx})}=-g_{3}^{(v_{xxz})}$. Theoretical justification for some of these symmetries is provided in appendix A.
An even larger number of terms are zero in the boundary conditions for the slip-no-slip surface, for the reason that some of the polynomial behaviours required at $Z\rightarrow\infty$ already have $u$ and its derivative $u_{z}$ (or $v$ and $v_{z}$) simultaneously vanish at the wall, and therefore satisfy the boundary conditions on both the slip and no-slip sections of the wall without any $Y$-dependence required. Such polynomial solutions do not entail any corrections and produce zero coefficients. In fact, as can be seen from figures 3 and 4, all the second-order velocity fields in the last column are either constant with $Y$ or of odd parity. Astoundingly enough, no second-order protrusion heights and only 8 third-order protrusion heights have non-zero values in the slip-no-slip case of table 1. Among these all coefficients involved in the $w(x,y,0,t)$ boundary condition (bottom row of (7.1)) are zero, for the trivial reason that the boundary condition $w(x,y,0,t)=0$ is exact.
8 Conclusions
As discussed in the introduction, the original concept of longitudinal and transverse protrusion heights, and the connection of their difference with the initial slope of the drag-reduction curve of riblets, have been extended by several authors over the years, both in order to further inform the design of drag-affecting surfaces by predicting their performance beyond the linear regime, and as equivalent slip-length boundary conditions to simplify numerical simulations of the same. This paper is an attempt to systematize such extensions by continuing the asymptotic expansion, of which the protrusion heights represent the first order, to higher orders, and in the while to discover the role of nonlinearities. The appropriate expansion parameter is unequivocally $\varepsilon=s^{+}$, the riblet period in wall units. The strength and at the same time the weakness of this approach is that it is a formal asymptotic expansion in the limit of $\varepsilon\rightarrow 0$, so on the positive side every term is strictly identified and there is no ambiguity as to what is negligible compared to what, on the negative side there is no guarantee that the approximation will be useful for practically interesting values of $\varepsilon$ (i.e., $s^{+}$). Nonetheless, there also is no guarantee for empirical approximations not based on a formal expansion, and at the very least the availability of approximations of multiple orders allows one to estimate a posteriori how close we are to reality by comparing successive orders.
The main operational results, at a first reading that skips the mathematics involved, are collected in the third-order equivalent boundary condition (7.1) and the table of its coefficients for five typical wall textures, table 1. We recall that the purpose of such an equivalent boundary condition is to replace the actual fine wall geometry, in the small-roughness limit where all of its dimensions tend to zero simultaneously, by a virtual flat wall where an approximate boundary condition is imposed. The advantage for numerical simulations, which otherwise become increasingly difficult as the texture becomes smaller and smaller in size, should be obvious.
Already deciding the appropriate, most general form of such an equivalent boundary condition was not evident before we started the present work. The practical answer is contained in the Cauchy expansion of an initial-value problem: a partial differential equation, or a system of those, can at least formally be specified as an initial-value problem in the wall-normal coordinate which for the Navier-Stokes equations involves six quantities, three components of the velocity vector and three components of the normal-stress vector, and their surface-parallel and time derivatives. The most general boundary condition that transforms this into an elliptic boundary-value problem involves three constraints among these six quantities. Most convenient is to write these three constraints in explicit form by assigning the three velocity components as functions of the three stress components and their wall-parallel derivatives, so that the solid-wall condition of zero velocity is naturally recovered when texture size $s^{+}$ tends to zero. This is what (7.1) does. One must note that, whereas wall-normal derivatives of stresses (equivalently replaced by $u_{z}$, $v_{z}$ and $p$) could in principle also appear in this formula, they can always be algebraically eliminated by using the differential equations themselves and their derivatives. Therefore a canonical form where wall-normal derivatives of $u_{z}$, $v_{z}$ and $p$ never appear can and should be used to eliminate ambiguities. This is notably also true of the wall-normal derivative $w_{z}$ of the wall-normal velocity, which can always be eliminated through the continuity equation and should not appear in the equivalent boundary condition at all.
The analysis proceeded through several intermediate problems, starting with the 2D Laplace and the 2D Stokes problem (which, not by accident, were the basic ingredients of the 1991 protrusion-height theory as well), and discovering step by step that these are also the building blocks of both 2D and 3D Navier-Stokes problems, when the wall geometry itself is 2D (as it is in classical riblets but not, for instance, in recent attempts at wavy riblets, or in general roughness). The limitation of the present article to 2D geometries, which anyway include a large number of practically interesting cases, is intentional as it gives the theory interesting distinctive features, but is not insurmountable. Effective boundary conditions for more general 3D textures were already introduced at first order by Sarkar & Prosperetti (1996) and by Luchini (2013), in the small-roughness limit, and at second order in the shallow-roughness limit by Kamrin et al. (2010), and could be generalized to higher orders in the future.
One outcome was to clarify, if there was any need, why only one longitudinal and one transverse protrusion heights were included in the (Luchini et al., 1991) original theory, despite the presence of two wall boundary conditions in the cross-flow problem. A formalization that specifies the most general relationship between the zeroth and first derivatives of the streamfunction $\psi$ on one side and its second and third derivatives on the other, through a matrix of four $a,b,c,d$ coefficients (4.3), shows that these coefficients are of different orders in the scale ratio $\varepsilon$, and $a$ alone has a role at first order. The $b,c,d$ coefficients only appear at higher orders, and were appropriately omitted by Luchini et al. (1991). They do play a role here, on the other hand.
Among the qualitative conclusions that can be drawn of the present analysis, even more interesting than what is included in the effective boundary condition (7.1) probably is what is not there. In particular, symmetry considerations forbid about half of possible couplings when riblets are left-right symmetric, because these coefficients would mix incompatible symmetries if present. We already mentioned that normal derivatives different from the basic shear components $u_{z}$, $v_{z}$ have no role in it, not so much because they cannot be there but because they need not be there; the Cauchy expansion process can always replace them with other quantities. In addition, undifferentiated pressure never appears (not so surprisingly because in incompressible flow pressure is defined up to an additive arbitrary constant), and this is another reason why the first-order equivalent boundary condition, in the first block of (7.1), only contains two protrusion heights.
At second order we see the appearance, still in the first line of (7.1), of an effect of the longitudinal and transverse pressure gradients on the respective wall-parallel velocity components, but also effects of shear-stress gradients $v_{yz}$ and $u_{xz}$ on the wall-normal velocity. All these effects have coefficients with the dimensions of an area, just because being of second order in the expansion they are proportional to the second power of the texture period $s^{2}$. The second derivative of wall normal velocity $w_{zz}$ need not appear because, from the continuity equation, it can always be replaced by $-u_{xz}-v_{yz}$. In addition, an unforeseen relationship was discovered between the effect of each pressure gradient on the corresponding velocity and the effect of the gradient of shear stress on normal velocity, in that their coefficients are necessarily equal, as evidenced in table 1 and justified in appendix A. The number of independent second-order protrusion heights for left-right symmetric riblets thus turns out to be just two. Two more second-order terms, only present for asymmetric riblets, involve cross effects of $v_{xz}$ upon longitudinal velocity $u$ and of $u_{xz}$ upon transverse velocity $v$, also with equal (aside from sign) coefficients. Two more coefficients, $h_{2}$ tying $u$ to $u_{yz}$, and $a_{2}$ tying $v$ to $v_{yz}$, will never be present for either symmetric or asymmetric riblets (see appendix B), despite these being the first that naturally turned up as “second-order protrusion heights” in the theory.
Concerning the slip-no-slip striped surface that is the simplest idealized model of a superhydrophobic surface, we found an unexpected important simplification in that all second-order effects are zero and the first non-zero corrections to the standard protrusion heights are of the third order. With hindsight this lack of second-order effects can be traced back to the lack of possible actions upon the normal velocity, because the impermeability condition is satisfied exactly, combined with the property above that states that the effects of pressure gradients have the same weights as the effects upon normal velocity.
But our perhaps most striking conclusion is that, amid the numerous third-order effects that are listed in table 1 and will not be repeated here, nonlinear terms are missing. Already that nonlinearities of the Navier-Stokes equations should not turn up before the third order was not obvious before the present formalization of the matched asymptotic expansion, but further when they do show up at third order they do so with a zero coefficient, not only for symmetric riblets (where these terms would be forbidden by symmetry), but also for any geometry as shown in appendix B. Therefore nonlinearity is repelled to fourth order (at least) in $s^{+}$, and linear equivalent boundary conditions can work unencumbered up to third order.
Acknowledgements. P. Luchini wishes to thank the Department of Mechanical Engineering of the University of Melbourne for a two-month invitation (March-April 2024) during which the core of this work was conceived.
Funding. D. Chung acknowledges support by the Air Force Office of Scientific Research (AFOSR) under award number FA2386-23-1-4071 (AOARD program manager: David Newell).
Appendix A Symmetry of the boundary-condition matrix
The Stokes equations satisfy a variational principle which gives them special symmetry properties. There is actually more than one version of this variational principle, but for the purposes of the biharmonic equation (4.2) its most convenient formulation uses the quantity
$J=\int_{\mathcal{D}}\frac{(\Delta_{2}\psi)^{2}}{2}\mathrm{d}S,$ (A.1)
$\mathcal{D}$ being the domain $0\leq Y\leq 1$, $F(Y)\leq Z\leq L$. Calculating the variation $\delta J$ gives
[…]
This is usually interpreted to say that, if boundary conditions are such as to nullify the boundary integral (for instance, if $\psi$ and $\partial\psi/\partial n$ are assigned on the whole boundary, so that their variations are zero), a solution of the biharmonic equation minimizes (A.1). However, more generally, this shows that if conditions are zero on part of the boundary and constant on the rest,
$\psi=\frac{\partial\psi}{\partial n}=0\text{ on }\partial\mathcal{D}_{1};\quad\frac{\partial\psi}{\partial n}=r_{1},\;\psi=r_{2}\text{ on }\partial\mathcal{D}_{2}$
one shall have
$\frac{\partial J}{\partial r_{1}}=\Delta_{2}\psi=s_{1},\quad\frac{\partial J}{\partial r_{2}}=-\frac{\partial\Delta_{2}\psi}{\partial n}=s_{2}.$
If now the solution of the equation is seen as a system with inputs $r_{j}$ and outputs $s_{i}$, $s_{i}$ will be linear functions of $r_{j}$, $s_{i}=A_{ij}r_{j}$. But
$A_{ij}=\frac{\partial^{2}J}{\partial r_{i}\partial r_{j}},$
and therefore $A_{ij}$ is a symmetric matrix, $A_{21}=A_{12}$. Similarly, if $s_{i}$ are assigned inputs and $r_{j}$ are outputs, the matrix tying one to the other, $A^{-1}$, will be symmetric. In the periodic flow problem with a riblet wall near $z=0$ and an opposite flat wall at distance $L\gg s$, the boundary integral vanishes on the riblet wall and the periodic boundaries, and $\partial\mathcal{D}_{2}$ is just the flat wall. In addition, on the flat wall conditions become asymptotically constant with $y$, and $\Delta_{2}\psi$ becomes just $\psi_{zz}$, as $L$ grows larger and larger. Therefore the relationship between $r_{1}=\partial\psi/\partial z,r_{2}=\psi,s_{1}=\psi_{zz}$, and $s_{2}=-\psi_{zzz}$ is symmetric. If one writes
$\left[\begin{matrix}\psi_{z}(y,L)\\ \psi(y,L)\end{matrix}\right]=\left[\begin{matrix}a&b\\ c&d\\ \end{matrix}\right]\left[\begin{matrix}\psi_{zz}(y,L)\\ -\psi_{zzz}(y,L)\end{matrix}\right],$
at distance $L$, the $[[a,b][c,d]]$ matrix will be symmetric. If this matrix is relayed back to $z=0$ using (4.7), it can be verified that it will still remain symmetric.
Another identity can be proved as follows. By integrating both sides of (6.11c) over $\mathcal{D}$ and applying the divergence theorem one gets
$-\int_{\mathcal{D}}\overline{\varPhi_{11}}\mathrm{d}Y\mathrm{d}Z=\int_{\mathcal{D}}\left(\overline{V_{23}}_{,Y}+\overline{W_{23}}_{,Z}\right)\mathrm{d}Y\mathrm{d}Z=\int_{0}^{1}\overline{W_{23}}(Y,L)\mathrm{d}Y;$
i.e., on account of (6.13), an alternate expression of $g_{2}$:
$g_{2}=\lim_{L\rightarrow\infty}\left(-\int_{\mathcal{D}}\overline{\varPhi_{11}}\mathrm{d}Y\mathrm{d}Z+h_{1}L+L^{2}/2\right).$ (A.2)
On the other hand, multiplying (3.9) by $\overline{\varPhi_{11}}$ and integrating gives
[…]
i.e., since $\overline{U_{22}}\sim Z^{2}/2+h_{2}^{(p_{x})}$, according to (3.10), and $\overline{\varPhi_{11}}\sim Z+h_{1}$, according to (2.13),
$\int_{\mathcal{D}}\overline{\varPhi_{11}}\mathrm{d}Y\mathrm{d}Z\sim(L+h_{1})L-(L^{2}/2+h_{2}^{(p_{x})})=L^{2}/2+h_{1}L-h_{2}^{(p_{x})}.$ (A.3)
Comparing (A.2) with (A.3) yields $g_{2}=h_{2}^{(p_{x})}$, consistently with what observed in table 1.
Four more symmetries were detected in table 1,
$f_{2}=-h_{2}^{(v_{xz})},\quad b_{3}=-c_{3},\quad h_{3}^{(p_{xy})}=-g_{3}^{(u_{xyz})},\quad f_{3}^{(p_{xx})}=-g_{3}^{(v_{xxz})},$
and will not be formally proved here. They have been, nonetheless, subjected to an additional numerical test using a riblet with smoothed edge, where numerical accuracy is much higher and allows them to be verified to several significant digits.
Appendix B Proof that some coefficients turn out to be zero for any geometry
It was observed in §7 that four coefficients of the effective boundary condition (7.1), among those that are expected to be zero for left-right symmetric wall geometries, are numerically zero even for an instance of asymmetric geometry. In fact these coefficients can be proved to be zero for any geometry, as follows.
B.1 $h_{2}$ (2.23)
The adjoint equation of (2.22) is
$\Delta_{2}\overline{\varPhi_{21}}^{\dagger}=0$ (B.1)
with boundary conditions $\overline{\varPhi_{21}}^{\dagger}[Y,F(Y)]=0$, $\overline{\varPhi_{21}}^{\dagger}_{,Z}(Y,\infty)=1$, and allows $h_{2}$ to be calculated from the r.h.s of (2.22) as
[…]
$\mathcal{D}$ being the domain $0\leq Y\leq 1$, $F(Y)\leq Z<\infty$.
But equation (B.1) and its boundary conditions are the same as (2.12), and therefore $\overline{\varPhi_{21}}^{\dagger}=\overline{\varPhi_{11}}$. It follows that
$h_{2}=-\frac{1}{2\pi}\int_{\mathcal{D}}\frac{\partial}{\partial Y}\left(\overline{\varPhi_{11}}^{2}\right)\mathrm{d}Y\mathrm{d}Z=0$
because $\overline{\varPhi_{11}}$ (and thus also $\overline{\varPhi_{11}}^{2}$) is periodic in $Y$ of period $1$, and the integral of the derivative of a periodic function is always zero.
B.2 $a_{2}$ (4.25)
The adjoint equation of (4.21) is
$\Delta_{2}\Delta_{2}\overline{\varPsi_{31}}^{\dagger}=0$ (B.2)
with boundary conditions
$\overline{\varPsi_{31}}^{\dagger}[Y,F(Y)]=\overline{\varPsi_{31}}^{\dagger}_{,N}[Y,F(Y)]=0,\quad\overline{\varPsi_{31}}^{\dagger}_{,ZZ}[Y,\infty]=1,\quad\overline{\varPsi_{31}}^{\dagger}_{,ZZZ}[Y,\infty]=0,$
and allows $a_{2}$ to be calculated from the r.h.s of (4.21) as
[…]
But equation (B.2) and its boundary conditions are the same as (4.14), and therefore $\overline{\varPsi_{31}}^{\dagger}=\overline{\varPsi_{21}}$. It follows that
[…]
for the same reasons as in §B.1.
B.3 $h_{3}^{(nl)}$ (6.17)
The adjoint equation of (6.16) is the same as (B.1), and therefore its solution $\overline{U_{34}}^{\dagger}$ also coincides with $\overline{\varPhi_{21}}^{\dagger}$ and with $\overline{\varPhi_{11}}$. It follows that (6.17) can be alternately calculated from the r.h.s of (6.16) as
$\begin{split}&h_{3}^{(nl)}=-\frac{1}{2\pi}\int_{\mathcal{D}}\varPhi_{1}\left(\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}_{,Z}\right)\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left[\overline{\varPsi_{21}}_{,Z}(\overline{\varPhi_{11}}^{2})_{Y}-\overline{\varPsi_{21}}_{,Y}(\overline{\varPhi_{11}}^{2})_{Z}\right]\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left(\overline{\varPsi_{21}}_{,Z}\overline{\varPhi_{11}}^{2}\right)_{Y}-\left(\overline{\varPsi_{21}}_{,Y}\overline{\varPhi_{11}}^{2}\right)_{Z}\mathrm{d}Y\mathrm{d}Z=0,\end{split}$
the last step coming from the divergence theorem and $\overline{\varPhi_{11}}^{2}[Y,F(Y)]=0$, $\overline{\varPsi_{21}}_{,Y}(Y,\infty)=0$.
B.4 $e_{3}^{(nl)}$ (5.15)
The adjoint equation of (5.13) is the same as (B.2), and therefore its solution $\overline{\varPsi_{44}}^{\dagger}$ also coincides with $\overline{\varPsi_{31}}^{\dagger}$ and with $\overline{\varPsi_{21}}$. It follows that the second of (5.15) can be alternately calculated from the r.h.s of (5.13) as
$\begin{split}&e_{3}^{(nl)}=-\frac{1}{2\pi}\int_{\mathcal{D}}\overline{\varPsi_{21}}\left(\overline{\varPsi_{21}}_{,Z}\Delta_{2}\overline{\varPsi_{21}}_{,Y}-\overline{\varPsi_{21}}_{,Y}\Delta_{2}\overline{\varPsi_{21}}_{,Z}\right)\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left[(\overline{\varPsi_{21}}^{2})_{Z}\Delta_{2}\overline{\varPsi_{21}}_{,Y}-(\overline{\varPsi_{21}}^{2})_{Y}\Delta_{2}\overline{\varPsi_{21}}_{,Z}\right]\mathrm{d}Y\mathrm{d}Z={}\\ &-\frac{1}{\pi}\int_{\mathcal{D}}\left[(\overline{\varPsi_{21}}^{2})_{Z}\Delta_{2}\overline{\varPsi_{21}}\right]_{Y}-\left[(\overline{\varPsi_{21}}^{2})_{Y}\Delta_{2}\overline{\varPsi_{21}}\right]_{Z}\mathrm{d}Y\mathrm{d}Z=0,\end{split}$
the last step coming from the divergence theorem and $(\overline{\varPsi_{21}}^{2})_{Y}[Y,F(Y)]=(\overline{\varPsi_{21}}^{2})_{Z}[Y,F(Y)]=0$, $\Delta_{2}\overline{\varPsi_{21}}(Y,\infty)=0$.